### v1

# Wasserstein Spatial Depth

Abstract
Modeling observations as random distributions embedded within Wasserstein spaces is becoming increasingly popular across scientific fields, as it captures the variability and geometric structure of the data more effectively. However, the distinct geometry and unique properties of Wasserstein space pose challenges to the application of conventional statistical tools, which are primarily designed for Euclidean spaces. Consequently, adapting and developing new methodologies for analysis within Wasserstein spaces has become essential. The space of distributions on $\mathbb{R}^{d}$ with $d>1$ is not linear, and “mimic” the geometry of a Riemannian manifold. In this paper, we extend the concept of statistical depth to distribution-valued data, introducing the notion of Wasserstein spatial depth. This new measure provides a way to rank and order distributions, enabling the development of order-based clustering techniques and inferential tools. We show that Wasserstein spatial depth (WSD) preserves critical properties of conventional statistical depths, notably, ranging within $[0,1]$, transformation invariance, vanishing at infinity, reaching a maximum at the geometric median, and continuity. Additionally, the population WSD has a straightforward plug-in estimator based on sampled empirical distributions. We establish the estimator’s consistency and asymptotic normality. Extensive simulation and real-data application showcase the practical efficacy of WSD.
62G05,
60F15,
62G20,
Depths,
Measure Transportation,
Outlier detection,
Wasserstein distance,
Wasserstein medians,
keywords:
[class=MSC]
keywords:
[…]
1 Introduction
Contemporary data collected in various disciplines is complex and multifaceted. Traditional statistical tools, which model data objects as samples from a Euclidean space or vector space, are inadequate to capture the variation and geometry of the data objects. Random objects lying in general metric spaces, including spaces of functions [45], Wasserstein spaces [31], and hyperbolic spaces [47], are gaining increasing favor in the scientific community. For instances, longitudinal images are treated as functions [45]; texts and media are modeled as distributions in modern AI training models [21, 10]; certain trees and graphs are embedded into hyperbolic spaces [9]. It is widely recognized that statistical efficiency can be gained by utilizing special properties of the above metric spaces [4].
In this paper, we focus on on modeling distribution-valued data objects within Wasserstein spaces. There are several advantages to model certain data objects as distributions or probability measures. Firstly, it captures the hierarchical variations in the data by simulating a two-stage data-generating process: initially sampling multiple distributions from a Wasserstein space, followed by drawing data points from each sampled distribution Secondly, it captures variations of the data along geodesics of the distribution space that are not straight lines as in the Euclidean setting and thus are closer to the observations. Thirdly, it often provides a low-dimensional embedding that effectively represents high-dimensional data, enabling better statistical inference without the curse of the dimension. Since the Wasserstein space has different structure and property from the Euclidean space, conventional analytic tools cannot apply to distribution-valued data objects. Therefore, new methods specifically designed for analyzing such data are essential.
There have been some efforts in this line of research, including but not limited to histogram regression [7], Wasserstein regression [12], geodesic PCA [5], template estimation [6], and Wasserstein clustering [48] or [18]. Despite the above development, there is limited effort in agnostic exploratory analysis for data objects in Wasserstein spaces [20, 16, 44, 19]. Still, exploratory analysis and descriptional statistics are critical to overview the properties of the data distribution before modeling. In particular, a notion of “ordering” for distribution-valued objects in Wasserstein spaces will be of fundamental utility. Besides exploratory analysis, it will also facilitate nonparametric methods for distribution-valued data.
Quantiles, ranks, and signs are pivotal tools of semiparametric and nonparametric statistics. Due to the lack of canonical ordering in multi-dimensional Euclidean space, quantile or rank based tools have been limited to one-dimensional data before the creation of statistical depths. The notion of statistical depths fills this gap, extending the notion of order to higher dimension. Given a distribution $P$ on $\mathbb{R}^{d}$, the depth of any data point $\mathbf{x}\in\mathbb{R}^{d}$ is a non-negative value that measures the “centrality” of $\mathbf{x}$ with respect to $P$. A larger value of depth indicates the data point is more central within the distribution, while data points with small depths are considered outliers or less typical within the distribution, worthy of investigation. Several different types of depths have been proposed, including Tukey depth [39, 33], simplicial depth [25], spatial depth [11, 41], Monge-Kantorovich depth [13, 22] and lens depth [26]. Via endowing multivariate data points with “center-outward” orderings, depths allow extension of order statistics, robust inferences and classifications to multivariate data [49, 30].
The concept of statistical depths has been extended to functional data, e.g., objects lying within a reproducing kernel Hilbert space (RKHS) [27, 32]. While it may be tempting to embed the Wasserstein space into a reproducing kernel Hilbert space (RKHS) and apply existing functional depth measures, this approach neglects the intricate geodesic structure of the Wasserstein space. There is no linear representation of the Wasserstein distance between distributions on $\mathbb{R}^{d}$ with $d>1$ [3]. Existing depths do not generalize well to nonlinear spaces. Besides the nonlinearity, the Wasserstein distance is computationally expensive even for empirical measures [35], which essentially rules out practical implementation of Tukey depth [39] and Monge-Kantorovich depth [13, 22]. The computational complexity of these two depths grows exponentially with the sample size.
In conclusion, conventional depth measures cannot be directly extended to Wasserstein spaces due to the unique properties and structure discussed above. This requires the development of a new notion of depth tailored specifically for Wasserstein spaces.
1.1 Contributions
In this paper we develop a new notion of depth to order or rank distributions. It is inspired by spatial depth, one of the simplest and most widely used notions of statistical depths. Recall that the spatial depth of a point $\mathbf{x}\in\mathbb{R}^{d}$ with respect to a probability measure $P$ over $\mathbb{R}^{d}$ is defined as
$\mathrm{SD}(\mathbf{x};P)=1-\left\|\mathbb{E}\left[\frac{\mathbf{X}-\mathbf{x}}{\|\mathbf{x}-\mathbf{X}\|}\right]\right\|,\quad\mathbf{X}\sim P.$
The spatial depth has been generalized to Hilbert spaces by following exactly the same definition [37, 41, 44]. However, the lack of linear structure of the Wasserstein space prevents an straightforward adaptation of the spatial depth. Nevertheless, the Wasserstein metric endows the space of probability measures with a structure of geodesic metric space (see [1]). For absolutely continuous probability measures $Q$ and $P$, the constant speed geodesic joining $Q$ and $P$ is given by the curve of probability measures
$[0,1]\ni\lambda\mapsto((1-\lambda){\rm I}+\lambda\,T_{Q,P})_{\#}Q,$
where ##\## denotes the push-forward operator (see Section 3.3 for its definition). The definition of spatial depth for manifold-valued data motivates us to define the depth of a probability measure $Q\in\mathcal{P}^{a.c}_{2}(\mathbb{R}^{d})$ (where $\mathcal{P}^{a.c}_{2}(\mathbb{R}^{d})$ is the set of absolutely continuous measures on $\mathbb{R}^{d}$ with finite second moments, see Section 3.3) with respect to a probability measure over the Wasserstein space $\mathbf{P}\in\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ (where $\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ is formally defined in Section 3.3) as
$\mathrm{SD}(Q;{\bf P}):=1-\left(\int{\left\|\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{\mathbf{x}-T_{Q,P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}\right]\right\|^{2}}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}}.$
Above, $T_{Q,P}$ is the optimal transport map from $Q$ to $P$ and $\mathcal{W}_{2}(P,Q)$ is the Wasserstein distance between $P$ and $Q$, both being formally defined in Section 3.3. We will show that $\mathrm{SD}(Q;\mathbf{P})$ satisfies the same properties as its Euclidean counterpart, namely transformation invariance, taking values in $[0,1]$, decreasing at infinity and attaining maximum at the median. Moreover, we show that $Q\mapsto\mathrm{SD}(Q;{\bf P})$ is continuous and ${\bf P}\mapsto\mathrm{SD}(Q;{\bf P})$ is continuous. Next, we will propose a finite-sample estimator under the so-called two-stage and one-stage sampling models. In the two-stage sampling model, such an estimator can be computed in polynomial time. Moreover, we will show that in both models, the estimator is consistent, meaning that it approximates the true population depth function as the sample size increases. We also prove asymptotic normality. Finally, we will provide numerical simulations for real and synthetic datasets.
In conclusion, Wasserstein spatial depth (WSD) serves as a valid measure for ordering objects within Wasserstein spaces, adhering to the axiomatic properties of depth [49] and being computationally feasible. This concept facilitates the extension of depth-based analytic tools to Wasserstein spaces, paving the way for future research.
1.2 Organization
General notations are provided in Section 2. The definition of WSD is given in Section 3 with illustrating examples in Section 4. In Section 5, we show that WSD shares the desirable properties of conventional statistical depths [49]. In Section 6, we tackle consistent estimation with asymptotic normality. In Section 7, we compare WSD to several depths in general metric spaces [20, 16, 44] adapted to Wasserstein spaces. We advocate WSD over the other depths in terms of computational feasibility, assumption flexibility, while possessing all desirable properties of a depth. In section 8, extensive numerical simulations are shown to demonstrate the empirical validity of WSD. Finally, in Section 9, we apply it to explore real-world data and make informative discoveries. All the proofs are provided in the appendix.
2 Notation
The space of Borel probability measures on a Polish space $(\mathcal{K},d)$ is denoted as $\mathcal{P}(\mathcal{K})$. For $P\in\mathcal{P}(\mathcal{K})$, its support is written ${\rm supp}(P)$. The space of Borel finite (signed) measures is denoted as $\mathcal{M}(\mathcal{K})$ and the space of finite (signed) measures with 00 mass as $\mathcal{M}_{0}(\mathcal{K})$, meaning that $h\in\mathcal{M}_{0}(\mathcal{K})$ if $h\in\mathcal{M}(\mathcal{K})$ and $h(\mathcal{K})=0$. The integral of a measurable function $f:\mathcal{K}\to\mathbb{R}$ with respect to $P\in\mathcal{P}(\mathcal{K})$ is denoted as
$\int f(\mathbf{x})\mathrm{d}P(\mathbf{x})=\int f\mathrm{d}P=P(f).$
Set $P\in\mathcal{P}(\mathcal{K})$ and $f:\mathcal{K}\to\mathbb{R}$ be measurable. Then
$\|f\|_{L^{2}(P)}:=\left(\int f^{2}\mathrm{d}P\right)^{1/2}$
denotes the $L^{2}(P)$-norm of $f$. The Hilbert space of measurable functions with finite $L^{2}(P)$-norm is denoted as $L^{2}(P)$ with inner product $\langle\cdot,\cdot\rangle_{L^{2}(P)}$. We say that a sequence $\{\mu_{n}\}_{n\in\mathbb{N}}\subset\mathcal{P}(\mathcal{K})$ converges weakly to $\mu\in\mathcal{P}(\mathbb{X})$ if
$\int f\,\mathrm{d}\mu_{n}\longrightarrow\int f\,\mathrm{d}\mu$
[…]
$d_{{\rm BL}}(\mu,\nu)\\ =\sup\bigg{\{}\int f(x)\mathrm{d}(\mu-\nu)(x):\ |f(x)|\leq 1\ {\rm and}\ |f(x)-f(y)|\leq d(x,y),\ \forall\,x,y\in\mathbb{X}\bigg{\}}.$
3 From Euclidean to Wasserstein spatial depth
In this section we define our notion of Wasserstein spatial depth. In Section 3.1 we recall the definition of Euclidean spatial depth and its main properties. In Section 3.2 we provide our interpretation of spatial depth in terms of geodesics, which allows its generalization to the Wasserstein space of measures (see Section 3.3). For readers interested in a more comprehensive understanding of the mathematical concepts discussed in Sections 3.2 and 3.3, we recommend consulting the monograph [1] for an in-depth exposition.
3.1 Euclidean spatial depth
In $\mathbb{R}^{d}$, for $d>1$, the spatial depth of a point $\mathbf{x}$ with respect to a random variable $\mathbf{X}\sim P$ is defined as
$\mathrm{SD}(\mathbf{x};P)=1-\left\|\mathbb{E}\left[\frac{\mathbf{X}-\mathbf{x}}{\|\mathbf{X}-\mathbf{x}\|}\right]\right\|.$
Throughout the paper, we use the convention $/0=0$ 0. The spatial depth shares the following properties with the univariate canonical depth function $2\min(F(x),1-F(-x))$. First, the statistical depth $\mathrm{SD}(\mathbf{x};P)$ belongs to the interval $[0,1]$. Second, the geometric median, defined as
${\bf m}_{\mathbf{X}}\in\operatorname*{arg\,min}_{\bf m}\mathbb{E}[\|\mathbf{X}-{\bf m}\|],$
satisfies $\mathrm{SD}({\bf m}_{\mathbf{X}};P)=1$. Third, as $\|\mathbf{x}\|\to\infty$, we have that $\mathrm{SD}(\mathbf{x};P)\to 0.$ Finally, for an isometric transformation $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$, it holds that
$\mathrm{SD}(T(\mathbf{x});T\#P)=\mathrm{SD}(\mathbf{x};P).$
3.2 Geodesic interpretations of the spatial depth
Let $(\mathcal{M},d)$ be a metric space. A curve $\{\gamma_{t}^{\mathbf{x}\to\mathbf{y}}\}_{t\in[0,1]}$ valued in $\mathcal{M}$ is a (constant speed) geodesic joining $\mathbf{x}\in\mathcal{M}$ to $\mathbf{y}\in\mathcal{M}$ if
$d(\gamma_{t}^{\mathbf{x}\to\mathbf{y}},\gamma_{s}^{\mathbf{x}\to\mathbf{y}})=(t-s)d(\mathbf{x},\mathbf{y}),\quad\text{for all }0\leq s\leq t\leq 1.$
The space $(\mathcal{M},d)$ is said to be geodesic if any two points are joined by at least one geodesic. The length of a curve $\{\gamma_{t}\}_{t\in[0,1]}$ with values in $\mathcal{M}$ (not necessarily a geodesic) is defined as $L(\gamma)=\int_{0}^{1}|\gamma_{t}^{\prime}|\mathrm{d}t$, where $|\gamma_{t}^{\prime}|=\lim_{s\to t}\frac{d(\gamma_{t},\gamma_{s})}{|t-s|}$. Assume now that $\mathcal{M}\subset\mathbb{R}^{d}$ is a Riemannian manifold with metric tensor $\{g_{\mathbf{x}}\}_{\mathbf{x}\in\mathcal{M}}$. Then it holds that
$L(\gamma)=\int_{0}^{1}\sqrt{g_{\gamma_{t}}(\partial_{t}\gamma_{t},\partial_{t}\gamma_{t})}\mathrm{d}t,$
where $\{\partial_{t}\gamma_{t}\}_{t\in[0,1]}$ denotes the velocity (standard time derivative) of the curve $\{\gamma_{t}\}_{t\in[0,1]}$.
In $\mathbb{R}^{d}$, a geodesic joining $\mathbf{x}$ and $\mathbf{y}$ is just the segment $\gamma_{t}^{\mathbf{x}\to\mathbf{y}}=(1-t)\mathbf{x}+t\mathbf{y}$, $t\in[0,1]$. Therefore, the spatial depth of $\mathbf{x}$ can be seen as the spatial depth of the velocities at time 00
$\mathrm{SD}(\mathbf{x};P)=1-\left\|\mathbb{E}\left[\frac{\partial_{t}|_{t=0}\gamma_{t}^{\mathbf{x}\to\mathbf{X}}}{\|\partial_{t}|_{t=0}\gamma_{t}^{\mathbf{x}\to\mathbf{X}}\|}\right]\right\|.$
This allows for the following Riemannian generalization of the spatial depth
$\mathrm{SD}(\mathbf{x};P)=1-\sqrt{g_{\mathbf{x}}\left(\mathbb{E}\left[\frac{\partial_{t}|_{t=0}\gamma_{t}^{\mathbf{x}\to\mathbf{X}}}{\|\partial_{t}|_{t=0}\gamma_{t}^{\mathbf{x}\to\mathbf{X}}\|}\right],\mathbb{E}\left[\frac{\partial_{t}|_{t=0}\gamma_{t}^{\mathbf{x}\to\mathbf{X}}}{\|\partial_{t}|_{t=0}\gamma_{t}^{\mathbf{x}\to\mathbf{X}}\|}\right]\right)}.$
3.3 Geodesic spatial depth over the space of measures
Let $\mathcal{P}_{p}(\mathbb{R}^{d})$ be the space of Borel probability measures over $\mathbb{R}^{d}$ with finite $p$th order moment. The optimal transport cost between two probability measures $P,Q\in\mathcal{P}_{p}(\mathbb{R}^{d})$ is defined as
$\mathcal{T}_{p}(P,Q)=\inf_{\pi\in\Pi(P,Q)}\frac{1}{2}\int\|\mathbf{x}-\mathbf{y}\|^{p}\mathrm{d}\pi(\mathbf{x},\mathbf{y}),$ (1)
where $\Pi(P,Q)\subset\mathcal{P}_{p}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ stands for the set of probability measures with marginals $P$ and $Q$, i.e., $(\mathbf{X},\mathbf{Y})\sim\pi\in\Pi(P,Q)$ if $\mathbf{X}\sim P$ and $\mathbf{Y}\sim Q$. For $p\geq 1$, the mapping $(P,Q)\mapsto\mathcal{W}_{p}(P,Q)=\left(\mathcal{T}_{p}(P,Q)\right)^{\frac{1}{p}}$ defines a distance over the space $\mathcal{P}_{p}(\mathbb{R}^{d})$ such that
$\mathcal{W}_{p}(P_{n},P)\to 0\quad\iff\quad P_{n}\xrightarrow{\mathcal{P}(\mathbb{R}^{d})}P\quad{\rm and}\quad\int\|\mathbf{x}\|^{p}\mathrm{d}P_{n}(\mathbf{x})\to\int\|\mathbf{x}\|^{p}\mathrm{d}P(\mathbf{x}).$
We focus now on the case $p=2$. We define $\mathcal{P}^{a.c}_{2}(\mathbb{R}^{d})$ as the subset of $\mathcal{P}_{2}(\mathbb{R}^{d})$ composed of absolutely continuous measures. If $P$ belongs to $\mathcal{P}^{a.c}_{2}(\mathbb{R}^{d})$, there exists a unique minimizer $\pi_{P,Q}$ of (1), for $p=2$. Moreover, there exists a unique gradient of a convex function $T_{P,Q}=\nabla\phi_{P,Q}$ such that $\pi_{P,Q}=({\rm I}\times\nabla\phi_{P,Q})_{\#}P$. The map $T_{P,Q}$ is called an optimal transport map. Here, for a probability measure $\mu$ and a Borel mapping $T$, $T_{\#}\mu$ denotes the push forward measure, which is the distribution of $T(\mathbf{X})$, for $\mathbf{X}\sim\mu$.
In [34], the author demonstrated that $\mathcal{W}_{2}$ serves as the natural metric for $\mathcal{P}_{2}(\mathbb{R}^{d})$, aimed at describing the long-term behavior of solutions to the porous medium equation. This metric also imparts a geodesic metric space structure to $\mathcal{P}_{2}(\mathbb{R}^{d})$.
It is natural in the following sense. If $\{\mathbf{X}_{t}\}_{t\in[0,1]}$ is a curve of random vectors with $\partial_{t}\mathbf{X}_{t}={\bf v}_{t}(\mathbf{X}_{0})$, then its associated curve of distributions $\{P_{t}\}_{t\in[0,1]}$ satisfies the so-called transport/continuity equation
$\partial_{t}P_{t}+{\rm div}({\bf v}_{t}P_{t})=0$ (2)
in an appropriate weak sense. The continuity equation is commonly used in fluid mechanics, where ${\bf v}_{t}$ represents the flow velocity vector field. However, given the curve $\{\mathbf{X}_{t}\}_{t\in[0,1]}$ there could exist several curves of velocity fields $\{{\bf v}_{t}\}_{t\in[0,1]}$ solving (2), i.e., generating the same flow. Among all of them, there exists only one belonging to
$\operatorname*{arg\,min}\left\{\int_{0}^{1}\|{\bf v}_{t}\|_{L^{2}(P_{t})}^{2}\mathrm{d}t:\quad\partial_{t}P_{t}+{\rm div}({\bf v}_{t}P_{t})=0\right\}.$ (3)
The tangent bundle of $(\mathcal{P}_{2}(\mathbb{R}^{d}),\mathcal{W}_{2})$ is
$\mathcal{T}_{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))=\overline{\{\nabla\phi:\quad\phi\in\mathcal{C}^{\infty}_{c}(\mathbb{R}^{d})\}}^{L^{2}(P)},\quad P\in\mathcal{P}_{2}(\mathbb{R}^{d})$
where $\mathcal{C}^{\infty}_{c}(\mathbb{R}^{d})$ denotes the set of infinitely differentiable functions with compact support. Above, $\overline{A}^{L^{2}(P)}$ denotes the closure of a subset $A$ in the Hilbert space $L^{2}(P)$. Given two probability measures $P$ and $Q$, a geodesic is any curve $\{\gamma_{t}^{P\to Q}\}_{t\in[0,1]}$ with endpoints $\gamma_{0}^{P\to Q}=P$ and $\gamma_{1}^{P\to Q}=Q$ with minimal velocity, i.e., any element of
$\operatorname*{arg\,min}\left\{\int_{0}^{1}\|{\bf v}_{t}\|_{L^{2}(\gamma_{t})}^{2}\mathrm{d}t:\quad\partial_{t}\gamma_{t}+{\rm div}({\bf v}_{t}\gamma_{t})=0,\ \gamma_{0}=P\ {\rm and}\ \gamma_{1}=Q\right\}.$ (4)
If $P$ belongs to $\mathcal{P}^{a.c}_{2}(\mathbb{R}^{d})$, there exists a unique geodesic given by the relation
$\gamma_{t}^{P\to Q}=((1-t){\rm I}+tT_{P,Q})_{\#}P.$
Its velocity field at $t=0$ is ${\bf v}_{0}^{P\to Q}=T_{P,Q}-{\rm I}$ and the Riemannian inner product in $\mathcal{T}_{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ is $\langle\cdot,\cdot\rangle_{L^{2}(P)}$. Therefore, the WSD of a probability measure $Q\in\mathcal{P}_{2}^{a.c.}(\mathbb{R}^{d})$ with respect to a a probability measure ${\bf P}$ over $\mathcal{P}_{2}(\mathbb{R}^{d})$ is defined as
$\mathrm{SD}(Q;{\bf P}):=1-\left\|\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{{\bf v}_{0}^{Q\to P}}{\|{\bf v}_{0}^{Q\to P}\|_{L^{2}(Q)}}\right]\right\|_{L^{2}(Q)}.$
Since, ${\bf v}_{0}^{Q\to P}=T_{Q\to P}-{\rm I}$ we get the following definition of spatial depth.
Definition 3.1.
The Wasserstein spatial depth of a distribution $Q\in\mathcal{P}^{a.c}_{2}(\mathbb{R}^{d})$ with respect to a distribution of distributions $\mathbf{P}\in\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ is defined as
$\mathrm{SD}(Q;{\bf P}):=1-\left(\int{\left\|\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{\mathbf{x}-T_{Q,P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}\right]\right\|^{2}}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}}.$
When $\mathbb{P}_{P\sim{\bf P}}(\mathcal{W}_{2}(P,Q)=0)\neq 0$, we set $\frac{\mathbf{x}-T_{Q,P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}=0$ for all $\mathbf{x}$ when $\mathcal{W}_{2}(P,Q)=0$.
Note that the definition of $\mathrm{SD}(Q;{\bf P})$ is focused on absolutely continuous distributions $Q$, while the distributions that ${\bf P}$ samples can be arbitrary (for instance, absolutely continuous, discrete, or a mixture of both). We also refer to the discussion in Section 10 on this point.
4 Examples
In this section we give several examples where the WSD can be computed explicitly.
4.1 Univariate case
In the case of univariate distributions, WSD reduces to quantile spatial depth. The univariate Wasserstein distance has a flat structure since there is an isometric homeomorphism between distributions and the corresponding CDF’s. Consequently, the Wasserstein distance between univariate distributions $P$ and $Q$ has a simple form
$\mathcal{W}_{2}^{2}(P,Q)=\int_{0}^{1}(F^{-1}_{P}(u)-F^{-1}_{Q}(u))^{2}\mathrm{d}u,$
where $F^{-1}_{P}(u)=\inf\{x\in\mathbb{R}:\ u\leq P((-\infty,x])\}$. Moreover, the univariate case is the unique case where the composition of optimal transport maps (here non-decreasing functions) is still an optimal transport map. Therefore, the spatial depth is just
$\mathrm{SD}(Q;{\bf P}):=1-\left(\int_{0}^{1}{\left(\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{F^{-1}_{P}(u)-F^{-1}_{Q}(u)}{\left(\int_{0}^{1}(F^{-1}_{P}(u)-F^{-1}_{Q}(u))^{2}\mathrm{d}\ell_{1}(u)\right)^{\frac{1}{2}}}\right]\right)^{2}}\mathrm{d}\ell_{1}(u)\right)^{\frac{1}{2}},$
which in short notation stands
$\mathrm{SD}(Q;{\bf P}):=1-\left\|{\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{F^{-1}_{P}-F^{-1}_{Q}}{\left\|F^{-1}_{P}(u)-F^{-1}_{Q}\right\|_{L^{2}([0,1])}}\right]}\right\|_{L^{2}([0,1])},$
which is the spatial depth of the quantile functions in the Hilbert space $L^{2}([0,1])$ (see [37, 41, 44]).
4.2 Location families
When $\mathbf{P}$ is supported on a location family, a set of shifted distributions indexed by the location parameter $\boldsymbol{\theta}$. In this case, $\mathbf{P}$ coincides with the distribution of the location parameter. And the WSD reduces to
$\mathrm{SD}(Q;{\bf P})=1-\left\|\mathbb{E}_{P\sim\mathbf{P}}\Bigg{(}\frac{\boldsymbol{\theta}_{P}-\boldsymbol{\theta}_{Q}}{\|\boldsymbol{\theta}_{P}-\boldsymbol{\theta}_{Q}\|}\Bigg{)}\right\|=1-\left\|\mathbb{E}_{\boldsymbol{\theta}}\Bigg{(}\frac{\boldsymbol{\theta}-\boldsymbol{\theta}_{Q}}{\|\boldsymbol{\theta}-\boldsymbol{\theta}_{Q}\|}\Bigg{)}\right\|,$ (5)
which is the Euclidean spatial depth of $\boldsymbol{\theta}_{Q}$ with respect to the distribution of $\boldsymbol{\theta}$. This also includes the Gaussian location family (see below).
4.3 Gaussian families
It is well known that the optimal transport problem between Gaussian probability measures admits a closed form (see [14]). In particular if $Q$ and $P$ are non degenerated Gaussian with means $\boldsymbol{\mu}_{Q}$ and $\boldsymbol{\mu}_{P}$ and (invertible) covariance matrices $\boldsymbol{\Sigma}_{Q}$ and $\boldsymbol{\Sigma}_{P}$, respectively, the optimal transport map $T_{Q,P}$ is
$\boldsymbol{\mu}_{P}+\boldsymbol{A}_{Q,P}(\mathbf{x}-\boldsymbol{\mu}_{Q})$
with
$\boldsymbol{A}_{Q,P}=\boldsymbol{\Sigma}_{Q}^{-\frac{1}{2}}\left(\boldsymbol{\Sigma}_{Q}^{\frac{1}{2}}\boldsymbol{\Sigma}_{P}\boldsymbol{\Sigma}_{Q}^{\frac{1}{2}}\right)^{\frac{1}{2}}\boldsymbol{\Sigma}_{Q}^{-\frac{1}{2}}.$
Therefore, if ${\rm supp}({\bf P})$ is a set of Gaussian probability measures and $Q$ is a non-degenerated Gaussian, then the WSD can be equivalently formulated as
$\mathrm{SD}(Q;{\bf P})=\\ 1-\left(\int{\left\|\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{\mathbf{x}-\boldsymbol{\mu}_{P}-\boldsymbol{A}_{Q,P}(\mathbf{x}-\boldsymbol{\mu}_{Q})}{\left(\|\boldsymbol{\mu}_{P}-\boldsymbol{\mu}_{Q}\|^{2}-{\rm Tr}\left(\boldsymbol{\Sigma}_{P}+\boldsymbol{\Sigma}_{Q}-2\left(\boldsymbol{\Sigma}^{\frac{1}{2}}_{P}\boldsymbol{\Sigma}_{Q}\boldsymbol{\Sigma}^{\frac{1}{2}}_{P}\right)^{\frac{1}{2}}\right)\right)^{\frac{1}{2}}}\right]\right\|^{2}}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}}.$
In the special case of a common $\boldsymbol{\Sigma}$ for all $P\in{\rm supp}(\mathbf{P})$, and when $\boldsymbol{\Sigma}_{Q}=\boldsymbol{\Sigma}$, the above formula reduces to
$\mathrm{SD}(Q;{\bf P})=1-\left\|\mathbb{E}_{P\sim\mathbf{P}}\Bigg{(}\frac{\boldsymbol{\mu}_{P}-\boldsymbol{\mu}_{Q}}{\|\boldsymbol{\mu}_{P}-\boldsymbol{\mu}_{Q}\|}\Bigg{)}\right\|,$
which is the Euclidean spatial depth function of $\boldsymbol{\mu}_{Q}$. When ${\rm supp}(\mathbf{P})=\frac{1}{n}\sum_{i=1}^{n}\delta_{P_{i}}$, we obtain
$\mathrm{SD}(Q;{\bf P}):=\\ 1-\left(\int{\left\|\frac{1}{n}\sum_{i=1}^{n}\left[\frac{\mathbf{x}-\boldsymbol{\mu}_{P_{i}}-\boldsymbol{A}_{Q,P_{i}}(\mathbf{x}-\boldsymbol{\mu}_{Q})}{\left(\|\boldsymbol{\mu}_{P_{i}}-\boldsymbol{\mu}_{Q}\|^{2}-{\rm Tr}\left(\boldsymbol{\Sigma}_{P_{i}}+\boldsymbol{\Sigma}_{Q}-2\left(\boldsymbol{\Sigma}^{\frac{1}{2}}_{P_{i}}\boldsymbol{\Sigma}_{Q}\boldsymbol{\Sigma}^{\frac{1}{2}}_{P_{i}}\right)^{\frac{1}{2}}\right)\right)^{\frac{1}{2}}}\right]\right\|^{2}}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}}.$
Furthermore, if $\boldsymbol{\Sigma}_{P_{i}}=\boldsymbol{\Sigma}_{Q}$ for all $i=1,\dots,n$, the WSD is
$1-\left(\int{\left\|\frac{1}{n}\sum_{i=1}^{n}\frac{\boldsymbol{\mu}_{P_{i}}-\boldsymbol{\mu}_{Q}}{\|\boldsymbol{\mu}_{P_{i}}-\boldsymbol{\mu}_{Q}\|}\right\|^{2}}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}}=1-\left\|\frac{1}{n}\sum_{i=1}^{n}\frac{\boldsymbol{\mu}_{P_{i}}-\boldsymbol{\mu}_{Q}}{\|\boldsymbol{\mu}_{P_{i}}-\boldsymbol{\mu}_{Q}\|}\right\|.$
5 Properties of Wasserstein spatial depth
Zuo and Serfling postulated in [49] the main four properties that a data depth should satisfy in Euclidean spaces. Those properties are affine invariance, meaning that the data depth function is invariant to affine transformations; center-outward monotonicity, meaning that the depth function decreases along rays arising from the deepest point; vanishing at infinity, meaning that the depth function tends to 00 as the distance to the deepest point tends to infinity; maximality at the center, meaning that for elliptic distributions, its geometric center is the unique deepest point. The Euclidean spatial depth satisfies some of these properties. In particular, it is invariant to isometric transformations, it vanishes at infinity and, if the spatial median is unique it is the unique maximizer of the spatial depth. As $\mathbb{R}^{d}$ is trivially embedded on $\mathcal{P}_{2}(\mathbb{R}^{d})$ by means of the mapping $\mathbf{x}\mapsto\delta_{\mathbf{x}}$, we cannot expect better properties for the Wasserstein space adaptation.
5.1 General properties
In this section we prove that the WSD shares the main properties of the Euclidean spatial depth, i.e., it belongs to the interval $[0,1]$, the geometric median is the deepest point, it decreases at infinite and it is transformation invariant.
Theorem 5.1.
Set $\mathbf{P}\in\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$. Then the following properties hold:
  1. 1.
    (Values in $[0,1]$.) $\mathrm{SD}(Q;{\bf P})\in[0,1]$ for all $Q\in\mathcal{P}^{a.c}_{2}(\mathbb{R}^{d})$.
  2. 2.
    (Transformation invariance.) Assume that $d\geq 2$. Then for any isometry $F:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathcal{P}_{2}(\mathbb{R}^{d})$, it holds that
    $\mathrm{SD}(F(Q);F_{\#}{\bf P})=\mathrm{SD}(Q;{\bf P}),\quad\text{for all $Q\in\mathcal{P}^{a.c}_{2}(\mathbb{R}^{d})$.}$
  3. 3.
    (Vanishing at infinity.) Let $\{Q_{n}\}_{n\in\mathbb{N}}\subset\mathcal{P}^{a.c}_{2}(\mathbb{R}^{d})$ be a sequence such that $\mathcal{W}_{2}(Q_{n},Q)\to+\infty$, for one $Q\in\mathcal{P}_{2}(\mathbb{R}^{d})$, then $\mathrm{SD}(Q_{n};{\bf P})\to 0.$
Recall from [3] that there are tree types of isometries in $(\mathcal{P}_{2}(\mathbb{R}^{d}),\mathcal{W}_{2})$. Let $F:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathcal{P}_{2}(\mathbb{R}^{d})$ be an isometry, i.e.,
$\mathcal{W}_{2}(F(P),F(Q))=\mathcal{W}_{2}(P,Q)\quad\text{for all }P,Q\in\mathcal{P}_{2}(\mathbb{R}^{d}).$
Then $F$ is called trivial if there exists an isometry $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ over $\mathbb{R}^{d}$ such that $F(P)=f_{\#}P$ for all $P\in\mathcal{P}_{2}(\mathbb{R}^{d})$; $F$ is said to preserve shapes if for all $P\in\mathcal{P}_{2}(\mathbb{R}^{d})$ there exists an isometry $f=f_{P}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $F(P)=f_{\#}P$; and if $F$ does not preserve shapes, is said to be exotic. An example of nontrivial isometry on $(\mathcal{P}_{2}(\mathbb{R}^{d}),\mathcal{W}_{2})$ that preserves shapes is given by the mapping $\Phi(\varphi):P\mapsto\Phi(\varphi)(P)$ where $\varphi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a linear isometry and $\Phi(\varphi)(P)$ is the law of the random variable
$\varphi(\mathbf{X}-\mathbb{E}[\mathbf{X}])+\mathbb{E}[\mathbf{X}],\quad{\rm for}\ \mathbf{X}\sim P.$
Theorems 1.1 and 1.2 in [3] proved that $(\mathcal{P}_{2}(\mathbb{R}^{d}),\mathcal{W}_{2})$ admits exotic isometries if and only if $d=1$, which is the reason for which the invariance of the WSD holds for $d\geq 2$.
5.2 Maximality at the center
The set of spatial medians of ${\bf P}$ is defined as
$\operatorname*{arg\,min}_{Q\in\mathcal{P}_{2}(\mathbb{R}^{d})}\mathbb{E}_{P\sim\mathbf{P}}[\mathcal{W}_{2}(P,Q)].$
The following result shows that, under some assumptions, the set of spatial medians which are absolutely continuous with respect to Lebesgue measure has maximum depth.
Theorem 5.2.
Set $\mathbf{P}\in\mathcal{P}(\mathcal{P}_{2}(\mathcal{K}))$ for a compact set $\mathcal{K}\subset\mathbb{R}^{d}$. Assume that $\mathbf{P}$ is supported on a finite set $P_{1},\dots,P_{n}$. Then any
$Q\in\mathcal{P}^{a.c}_{2}(\mathcal{K})\cap\operatorname*{arg\,min}_{Q^{\prime}\in\mathcal{P}(\mathcal{K})}\mathbb{E}_{P\sim\mathbf{P}}[\mathcal{W}_{2}(P,Q^{\prime})]$
such that $Q\neq P_{i}$ for all $i=1,\dots n$ satisfies $\mathrm{SD}(Q;{\bf P})=1$.
Remark 5.3.
We do not know if the set of spatial medians which are absolutely continuous with respect to Lebesgue measure is nonempty. It is known that, under the setting of Theorem 5.2, if we assume that $P_{i}\in\mathcal{P}^{a.c}_{2}(\mathcal{K})$, the set of geometric means (or barycenters) is a singleton and its unique element belongs to $\mathcal{P}^{a.c}_{2}(\mathcal{K})$ (see [50]). However, the proof of [50], based on a fixed point argument which exploits the strict convexity of the squared Wasserstein distance, does not apply to our setting.
5.3 Continuity
In this section we investigate some topological properties of the WSD. We analyze separately the function $Q\mapsto\mathrm{SD}(Q;{\bf P})$ and ${\bf P}\mapsto\mathrm{SD}(Q;{\bf P})$. The following result shows that that the function $\mathcal{P}_{2}^{a.c.}(\mathbb{R}^{d})\ni Q\mapsto\mathrm{SD}(Q;{\bf P})$ is continuous.
Theorem 5.4.
Let $\mathbf{P}\in\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ be atomless and $\{Q_{n}\}_{n\in\mathbb{N}}\subset\mathcal{P}_{2}^{a.c.}(\mathbb{R}^{d})$ be a sequence such that $\mathcal{W}_{2}(Q_{n},Q)\to 0$ for some $Q\in\mathcal{P}_{2}^{a.c.}(\mathbb{R}^{d})$. Then
$\lim_{n\to\infty}\mathrm{SD}(Q_{n};{\bf P})=\mathrm{SD}(Q;{\bf P}).$
Next we show the continuity of ${\bf P}\mapsto\mathrm{SD}(Q;{\bf P})$ for fixed $Q$. As in intermediate step we need to show that that for each $Q\in\mathcal{P}_{2}^{a.c.}(\mathbb{R}^{d})$ the function
$T^{Q}:\mathcal{P}_{2}(\mathbb{R}^{d})\ni P\mapsto T_{Q,P}\in L^{2}(Q)$
is continuous.
Lemma 5.5 (Continuity of $T^{Q}$).
Set $Q\in\mathcal{P}_{2}^{a.c}(\mathbb{R}^{d})$. Let $\{P_{n}\}_{n}\subset\mathcal{P}_{2}(\mathbb{R}^{d})$ be a sequence of probability measures such that $\mathcal{W}_{2}(P_{n},P)\to 0$ for some $P\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Then
$\|T_{Q,P_{n}}-T_{Q,P}\|_{L^{2}(Q)}\to 0.$
In words, $T^{Q}:\mathcal{P}_{2}(\mathbb{R}^{d})\to L^{2}(Q)$ is continuous.
Fix $Q\in\mathcal{P}_{2}^{a.c}(\mathbb{R}^{d})$. Lemma 5.5 implies that the function
$\mathcal{P}_{2}(\mathbb{R}^{d})\ni P\mapsto\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}\in L^{2}(Q)$
is continuous around all $P\neq Q$. This observation enables to derive the continuity of the function $\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))\ni{\bf P}\mapsto\mathrm{SD}(Q;{\bf P})$ around atomless probability measures.
Theorem 5.6.
Let $\mathbf{P}\in\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ be atomless and let $Q\in\mathcal{P}_{2}^{a.c}(\mathbb{R}^{d})$. Then
$\lim_{n\to\infty}\mathrm{SD}(Q;{\bf P}_{n})=\mathrm{SD}(Q;{\bf P})$
for every sequence $\{{\bf P}_{n}\}_{n\in\mathbb{N}}\subset\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ such that ${\bf P}_{n}\rightarrow{\bf P}$ weakly in $\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$.
6 Consistent estimation
In practice, we only observe sample datasets instead of knowing the true ${\bf P}$ or even the true $P_{1},P_{2},\ldots,P_{n}\sim{\bf P}$. Two common scenarios in the literature of distributional data learning [2, 29, 38] will be considered, namely, one-stage sampling model and two-stage sampling model. One-stage sampling model assumes the observation of an i.i.d. sample $P_{1},\dots,P_{n}$ of ${\bf P}$. Two-stage sampling model assumes the observation of a data array
$\left(\begin{array}[]{ccc}\mathbf{X}_{1,1}&\dots&\mathbf{X}_{1,m}\\ \vdots&\vdots&\vdots\\ \mathbf{X}_{n,1}&\dots&\mathbf{X}_{n,m}\end{array}\right),$ (6)
where $\mathbf{X}_{i,1},\dots,\mathbf{X}_{i,m}\in\mathbb{R}^{d}$ is an i.i.d. sample from the distribution $P_{i}$ for each $i=1,\dots,n$, and $P_{1},\dots,P_{n}$ are i.i.d. drawn from ${\bf P}$. The difference between the two models is that the sampled distributions $P_{1},\ldots,P_{n}$ are known in one-stage sampling model, but unknown and to be estimated by the empirical distributions in two-stage sampling model.
In each scenario, we give the empirical counterpart to the population WSD in Definition 3.1. We also establish a point-wise central limit theorem for the empirical WSD under the one-stage sampling model and a consistency result for the two-stage sampling model.
6.1 One-stage sampling
We describe the asymptotic behavior of the empirical WSD
$\mathrm{SD}(Q;{\bf P}_{n}):=1-\left(\int{\left\|\frac{1}{n}\sum_{i=1}^{n}\left[\frac{\mathbf{x}-T_{Q,P_{i}}(\mathbf{x})}{\mathcal{W}_{2}(Q,P_{i})}\right]\right\|^{2}}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}},$
which is built upon the empirical measure ${\bf P}_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{P_{i}}$. The WSD is associated with the spatial distribution process
$S_{{\bf P}}:\mathcal{P}^{a.c}_{2}(\mathbb{R}^{d})\ni Q\mapsto S_{{\bf P},Q}=\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}\right]\in L^{2}(Q).$
The representation
$S_{{\bf P},Q}=\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{I-T_{Q,P}}{\|I-T_{Q,P}\|_{L^{2}(Q)}}\right]$
allows to use standard techniques to obtain the point-wise strong law of large numbers and a central limit theorem for the empirical spatial distribution process $\left(S_{{\bf P}_{n},Q}-S_{{\bf P},Q}\right)\in L^{2}(Q),$ after showing that the random function $T_{Q,P}$ in $L^{2}(Q)$ is tight if ${P}\sim{\bf P}$ for a tight ${\bf P}$. In other words, we need that for each $Q\in\mathcal{P}_{2}^{a.c.}(\mathbb{R}^{d})$ the function
$T^{Q}:\mathcal{P}_{2}(\mathbb{R}^{d})\ni P\mapsto T_{Q,P}\in L^{2}(Q)$
pushes forward tight probability measures over $\mathcal{P}_{2}(\mathbb{R}^{d})$ to tight probabilities in $L^{2}(Q)$. Recall that a probability measure $\mu\in\mathcal{P}(\mathcal{X})$ over a separable topological space $(\mathcal{X},d)$ is said to be tight if for every $\epsilon>0$ there exists a compact (in the metric topology) set $K$ such that $\mu(K)\geq 1-\epsilon.$ A random variable is tight if its distribution is tight. Therefore, Lemma 5.5 implies that if $P\sim{\bf P}$ is tight in $\mathcal{P}_{2}(\mathbb{R}^{d})$, then $T_{Q,P}$ is tight in $L^{2}(Q)$. As a consequence, if ${\bf P}\in\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$, then $\left\{\frac{I-T_{Q,P_{i}}}{\|I-T_{Q,P_{i}}\|_{L^{2}(Q)}}\right\}_{i=1}^{n}$ is an i.d.d. sequence of tight random elements in the separable Hilbert space $L^{2}(Q)$, with finite second order moments. The strong law of large numbers and the central limit theorem in separable Hilbert spaces (cf. [24, Corollary 10.9]) yield the following result. Note that a random element $Z$ of a Hilbert space $\mathcal{H}$ is defined to be Gaussian when $h(Z)$ follows a (univariate) Gaussian distribution for all linear continuous mappings $h:\mathcal{H}\to\mathbb{R}$.
Theorem 6.1.
Set $\mathbf{P}\in\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$, $Q\in\mathcal{P}^{a.c}_{2}(\mathbb{R}^{d})$. Then
$\|S_{{\bf P}_{n},Q}-S_{{\bf P},Q}\|_{L^{2}(Q)}\xrightarrow{a.s}0$
and
$\sqrt{n}(S_{{\bf P}_{n},Q}-S_{{\bf P},Q})\xrightarrow{\mathcal{P}(L^{2}(Q))}\mathbb{G}_{{\bf P},Q},$
for some centered Gaussian element $\mathbb{G}_{{\bf P},Q}\in L^{2}(Q)$. As a consequence, it holds that
$\mathrm{SD}(Q;{\bf P}_{n})\xrightarrow{a.s}\mathrm{SD}(Q;{\bf P})$
and, if $\mathrm{SD}(Q;{\bf P})<1$, also
$\sqrt{n}(\mathrm{SD}(Q;{\bf P}_{n})-\mathrm{SD}(Q;{\bf P}))\xrightarrow{\mathcal{P}(\mathbb{R})}\frac{\langle\mathbb{G}_{{\bf P},Q},S_{{\bf P},Q}\rangle_{L^{2}(Q)}}{\mathrm{SD}(Q;{\bf P})-1}.$
The last two statements of Theorem 6.1 are a mere application of the delta method.
6.2 Two-stage sampling
Now we deal with the scenario where only the data (6) array is available. Recall that, in this case, the i.i.d. samples $P_{1},\dots,P_{n}$ of ${\bf P}$ are no longer observed but a sample $\{\mathbf{X}_{i,j}\}_{i,j=1}^{n,m}$ is, where $\mathbf{X}_{i,j}\sim P_{i}$ for $j\in\{1,\dots m\}$ and each $i\in\{1,\dots n\}$ We denote
${\bf P}_{n,m}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{P_{i,m}},\quad\text{with}\quad P_{i,m}:=\frac{1}{m}\sum_{j=1}^{m}\delta_{\mathbf{X}_{i,j}}\quad\text{for each}\quad i\in\{1,\dots n\}.$ (7)
Correspondingly, the empirical WSD is formulated as
$\mathrm{SD}(Q_{m};{\bf P}_{n,m})=1-\sqrt{\frac{1}{m}\sum_{j=1}^{m}\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\frac{X_{q,j}-T_{Q_{m},P_{i,m}}(X_{q,j})}{\mathcal{W}_{2}(Q_{m},P_{i,m})}\bigg{\|}^{2}},$
where $Q_{m}=\frac{1}{m}\sum_{j=1}^{m}\delta_{X_{q,j}}$ with $X_{q,1},\ldots,X_{q,m}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}Q$, and the convention $\frac{\boldsymbol{0}}{\boldsymbol{0}}=\boldsymbol{0}$ remains. Now we show that, as $n,m\to\infty$, ${\bf P}_{n,m}$ converges in probability in $\mathcal{P}(\mathcal{P}_{p}(\mathbb{R}^{d}))$ for all $p\geq 1$. We endow $\mathcal{P}(\mathcal{P}_{p}(\mathbb{R}^{d}))$ with the metric
$d_{{\rm BL}^{(p)}}({\bf P},{\bf Q})\\ =\sup\bigg{\{}\int f(P)\mathrm{d}({\bf P}-{\bf Q})(P):\ |f(P)|\leq 1\ {\rm and}\ |f(P)-f(Q)|\leq\mathcal{W}_{p}(P,Q),\ \forall\,P,Q\in\mathcal{P}_{p}(\mathbb{R}^{d})\bigg{\}}.$
Lemma 6.2.
Let $\{{\bf P}_{n,m}\}_{n\in\mathbb{N}}$ be as in (7) where $m=m(n)$ is such that $m\to\infty$ as $n\to\infty$. Assume that ${\bf P}\in\mathcal{P}(\mathcal{P}_{p}(\mathbb{R}^{d}))$ for $p\geq 1$. Then
$\mathbb{E}[d_{{\rm BL}^{(p)}}({\bf P}_{n,m},{\bf P})]\longrightarrow 0\quad\text{as }n\to\infty.$
A combination of Lemma 6.2, Theorem 5.6 and the continuous mapping theorem yields the following consistency result for the two-staged sample estimator.
Theorem 6.3.
Set $\mathbf{P}\in\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ be atomless. Let $\{{\bf P}_{n,m}\}_{n\in\mathbb{N}}$ be as in (7) where $m=m(n)$ is such that $m\to\infty$ as $n\to\infty$. Then, for every $Q\in\mathcal{P}_{2}(\mathbb{R}^{d})$,
$\mathrm{SD}(Q;{\bf P}_{n,m})\overset{\mathbb{P}}{\longrightarrow}\mathrm{SD}(Q;{\bf P})\quad\text{as }n\to\infty.$
7 Comparison with other possible depth notions
In the field of nonparametric statistics, the concept of depths presents significant challenges when attempting generalization to non-Euclidean spaces, a topic that has garnered considerable attention in advanced statistical research. Within the confines of linear functional spaces, such as Banach or Hilbert spaces, the application of Euclidean methodologies remains largely successful, attributed primarily to their inherent vectorial structures. Contrastingly, the landscape becomes markedly more complex when venturing into the domain of infinite-dimensional spaces devoid of a vectorial framework.
The statistical literature identifies a mere trio of propositions capable of addressing this complexity: lens depth, Tukey depth, and a novel approach of metric spatial depth, different from our proposal. Here we delve into a meticulous exploration of these methodologies, with a particular emphasis on their adaptability to Wasserstein space framework.
We shall demonstrate that these methodologies do not possess all the favorable theoretical and computational properties that we have established for the WSD. The WSD is thus most beneficial in broad, complex statistical contexts, thereby yielding a significant advancement in the field of machine learning and statistical analysis.
7.1 Tukey depth
The next definition is a natural adaptation of the metric Tukey (or halfspace) depth proposed by Dai and Lopez-Pintado [46] to the Wasserstein space.
Definition 7.1 (Adapted from [46]).
The Wasserstein halfspace depth of a distribution $Q\in\mathcal{P}_{2}(\mathbb{R}^{d})$ with respect to a probability measure $\mathbf{P}\in\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ is the value
$\mathrm{HSD}(Q;{\bf P})=\inf_{\begin{subarray}{c}P_{1},P_{2}\in\mathcal{P}_{2}(\mathbb{R}^{d})\\ \mathcal{W}_{2}(Q,P_{1})\leq\mathcal{W}_{2}(Q,P_{2})\end{subarray}}\mathcal{P}_{P\sim\mathbf{P}}(\mathcal{W}_{2}(P,P_{1})\leq\mathcal{W}_{2}(P,P_{2})).$
According to [46], the Wasserstein halphspace depth is transformation invariant and vanishes at infinity. Moreover, center-outward monotonicity (the function $t\mapsto\mathrm{HSD}(\gamma(t);{\bf P})$ is monotone decreasing for any geodesic $\gamma(t)$ with $\mathrm{HSD}(\gamma(0);{\bf P})=1/2$) holds if for any constant speed geodesic $\gamma$ of $(\mathcal{P}_{2}(\mathbb{R}^{d}),\mathcal{W}_{2})$, the following geometric condition holds:
if there there exists t[0,1] such that $\text{if there there exists $t\in[0,1]$ such that }\mathcal{W}_{2}^{2}(\gamma(t),P)\leq\mathcal{W}_{2}^{2}(\gamma(t),Q)\\ \implies\left(\mathcal{W}_{2}^{2}(\gamma(0),P)\leq\mathcal{W}_{2}^{2}(\gamma(0),Q)\right)\ {\rm or}\ \left(\mathcal{W}_{2}^{2}(\gamma(1),P)\leq\mathcal{W}_{2}^{2}(\gamma(1),Q)\right).$ (8)
Recall that a constant speed geodesic in a metric space $(\mathcal{M},d)$ is a curve $\gamma:[0,1]\to\mathcal{M}$ such that $d(\gamma(t),\gamma(s))=|t-s|d(\gamma(0),\gamma(1))$ for all $s,t\in[0,1]$. In $(\mathcal{P}_{2}(\mathbb{R}^{d}),\mathcal{W}_{2})$, a constant speed geodesic corresponds to interpolations obtained from optimal transport plans [1, p. 158]. More precisely, any constant speed geodesic connecting two absolutely continuous distributions $P_{1}$ and $P_{2}$ is of the form $\gamma(t)=((1-t)I+tT)_{\#}P_{1}$, where $T$ is the unique optimal map pushing $P_{1}$ to $P_{2}$.
Center-outward monotonicity is widely regarded as a favorable attribute within the scope of statistical depth measures. However, it is an attribute not typically anticipated in the context of spatial depths, particularly within Euclidean spaces. Notably, neither the transport-based depth nor the lens depth exhibit this property. Moreover, for the Wasserstein halphspace depth it is not clear if the geometric condition (8), and a fortiori the center outward monotonicity, holds in general.
Despite the ostensibly advantageous traits of Tukey depths, they are encumbered by significant computational demands, particularly evident as the dimensionality of the data increases. This computational intensity escalates to the point of impracticality for exact calculations in dimensions exceeding five, already in the Euclidean case. Within the confines of Wasserstein spaces, which are characterized by infinite dimensions, approximating Tukey depths poses a substantial challenge, much more so than for the WSD.
7.2 Lens depth
Let us now turn our attention to the adaptation of the metric lens depth, presented by Geenens, Nieto-Reyes and Francisci in [20], to the Wasserstein space.
Definition 7.2 (Adapted from [20]).
The Wasserstein lens depth of a distribution $Q\in\mathcal{P}_{2}(\mathbb{R}^{d})$ with respect to ${\bf P}\in\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ is defined as
$\mathrm{LD}(Q;{\bf P})=\mathcal{P}_{(P^{\prime},P)\sim{\bf P}\otimes{\bf P}}[\mathcal{W}_{2}(P,P^{\prime})\geq\max(\mathcal{W}_{2}(P,Q),\mathcal{W}_{2}(P^{\prime},Q))].$
The Wasserstein lens depth is transformation invariant and vanishes at infinity. The two-stage plug-in estimator of $\mathrm{LD}(Q;{\bf P})$ can be computed exactly for a discrete distribution $Q$ within polynomial time. Nevertheless, as indicated in [20], the lens depth fails to exhibit center-outward monotonicity in the linear case. Similarly, this property would possibly be absent in Wasserstein spaces.
7.3 Metric spatial Wasserstein depth
Virta [44] gave a definition of spatial depth for general metric spaces that does not agree with our definition in the particular case of Wasserstein space. To avoid confusion in terminology, the proposal from [44] will be referred to as metric spatial Wasserstein depth.
Definition 7.3 (Adapted from [44]).
The metric spatial Wasserstein depth of $Q\in\mathcal{P}_{2}(\mathbb{R}^{d})$ with respect to ${\bf P}\in\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ is defined as
$\mathrm{MSD}(Q;{\bf P})=1-\frac{1}{2}\mathbb{E}_{(P^{\prime},P)\sim{\bf P}\otimes{\bf P}}\left[\frac{\mathcal{W}_{2}^{2}(P,Q)+\mathcal{W}_{2}^{2}(P^{\prime},Q)-\mathcal{W}_{2}^{2}(P,P^{\prime})}{\mathcal{W}_{2}(Q,P)\mathcal{W}_{2}(Q,P^{\prime})}\right].$
The function $Q\mapsto\mathrm{MSD}(Q;{\bf P})$ takes values in the interval $[0,2]$. It is transformation invariant and vanishing at infinity. The metric spatial depth presents a remarkably viable and effective solution that is widely applicable to general metric spaces. Nevertheless, when specialized to the Wasserstein space, it falls short of fulfilling all the desirable properties that we have established for the WSD. In particular, also pointed out in [44], the question of the inclusion of spatial medians within the set of deepest points in terms of the metric spatial depth remains overall open. Note that taking a directional derivative of $\mathrm{MSD}(Q;{\bf P})$ with respect to $Q$, in the aim of studying deepest points, does not seem particularly fruitful. This leads us to conjecture that, in general, spatial medians have no relation to the maximizers of $\mathrm{MSD}(Q;{\bf P})$. In contrast, our Theorem 5.2 establishes that spatial medians maximize the WSD, in more general situations.
Another natural question remaining overall open in [44] is whether the maximal possible value $2$ for the metric spatial depth can be reached. In particular Theorem 3 there states that this value $2$ is attained by any non-atomic $Q$ such that, given two independent realizations $P_{1}$ and $P_{2}$ from P, $Q$ always falls between these two points on a geodesic going through all three. As noted in [44], this condition is very strict and only satisfied in arguably pathological cases.
8 Numerical simulations
In this section, we carry out extensive numerical simulations to validate our notion of Wasserstein spatial depth and support its theoretical properties and practical utility. Specifically, we confirm the consistency of the empirical WSD, evaluate its effectiveness in outlier detection, and examine its relationship with conventional spatial depth in certain cases. Throughout this section, we use the R package transport to compute all the Wasserstein distances and optimal transport maps from data clouds. Based on the two stage sampling model in (6), the empirical WSDs are calculated via the formula below. For $Q_{m}=\frac{1}{m}\sum_{j=1}^{m}\delta_{X_{q,j}}$ with $X_{q,1},\ldots,X_{q,m}\stackrel{{\scriptstyle\textit{i.i.d. }}}{{\sim}}Q$,
$\mathrm{SD}(Q_{m};{\bf P}_{n,m})=1-\sqrt{\frac{1}{m}\sum_{j=1}^{m}\bigg{\|}\frac{1}{n}\sum_{i\neq q}\frac{X_{q,j}-T_{Q_{m},P_{i,m}}(X_{q,j})}{\mathcal{W}_{2}(Q_{m},P_{i,m})}\bigg{\|}^{2}},$ (9)
where $Q_{m}$ could be outside (with the convention $q=n+1$) or within (with the convention $q\in\{1,\ldots,n\}$) the sampled distributions $\{P_{1,m},\ldots,P_{n,m}\}$, and where we recall the convention $/0=0$ 0. The code for all simulations is publicly available at https://github.com/YishaYao/Wasserstein-Spatial-Depth/tree/main.
Since computing the OT map between any pair of empirical distributions costs $O(m^{2})$ and once the OT map between a pair of empirical distributions is available, the corresponding Wasserstein distance immediately follows with almost zero extra cost, the computational complexity of $SD(Q_{m},\boldsymbol{P}_{n,m})$ is of order $O(nm^{2})$.
8.1 Consistency of the empirical Wasserstein spatial depth
The simulation results below support the theoretical results in Section 6. That is, the empirical WSD, formulated in (9), is close to the theoretical value $\mathrm{SD}(Q;{\bf P})$ in Definition 3.1. Hence, the WSD can be inferred accurately from sample data and has practical value. Four cases are considered and described below.
  • Case 1: ${\bf P}$ is supported on a family of exponential distributions indexed by the rate parameter $\lambda$ which follows a $\mathrm{Beta}(2,2)$ distribution. The theoretical WSD of the exponential distribution with rate parameter $\lambda_{Q}\in(0,1]$, denoted as $\mathrm{exp}(\lambda_{Q})$, with respect to ${\bf P}$ is
    $\mathrm{SD}(Q;{\bf P})$ $=1-\sqrt{\int_{0}^{\infty}\Big{(}\mathbb{E}_{\lambda\sim\mathrm{Beta}(2,2)}\frac{x-(\lambda_{Q}/\lambda)x}{\mathcal{W}_{2}(F_{\lambda_{Q}},F_{\lambda})}\Big{)}^{2}\lambda_{Q}e^{-\lambda_{Q}x}\mathrm{d}x}$
    $=1-\sqrt{\int_{0}^{\infty}\frac{\lambda_{Q}^{2}x^{2}}{2}\Big{(}\mathbb{E}_{\lambda\sim\mathrm{Beta}(2,2)}\frac{1/\lambda_{Q}-1/\lambda}{|1/\lambda_{Q}-1/\lambda|}\Big{)}^{2}\lambda_{Q}e^{-\lambda_{Q}x}\mathrm{d}x}$
    $=1-\sqrt{\int_{0}^{\infty}\frac{\lambda_{Q}^{2}}{2}\Big{(}4\lambda_{Q}^{3}-6\lambda_{Q}^{2}+1\Big{)}^{2}x^{2}\lambda_{Q}e^{-\lambda_{Q}x}\mathrm{d}x}$
    $=1-\Big{|}1+4\lambda_{Q}^{3}-6\lambda_{Q}^{2}\Big{|},$
    where $F_{\lambda}$ is the CDF of the exponential distribution with rate parameter $\lambda$, the optimal map from $\mathrm{exp}(\lambda_{Q})$ to $\mathrm{exp}(\lambda)$ is
    $T_{\lambda_{Q},\lambda}(x)=F_{\lambda}^{-1}\circ F_{\lambda_{Q}}(x)=\frac{\lambda_{Q}x}{\lambda},$
    and $\mathcal{W}_{2}(F_{\lambda_{Q}},F_{\lambda})$ is derived by
    $\mathcal{W}_{2}(F_{\lambda_{Q}},F_{\lambda})=\sqrt{\int_{0}^{\infty}\Big{(}x-(\lambda_{Q}/\lambda)x\Big{)}^{2}\lambda_{Q}e^{-\lambda_{Q}x}\mathrm{d}x}=\sqrt{2}\Big{|}\frac{1}{\lambda_{Q}}-\frac{1}{\lambda}\Big{|}.$
    Note that the WSD is equal to $1$ (maximal) for $\lambda_{Q}=1/2$ which is the mean of the $\mathrm{Beta}(2,2)$ distribution.
  • Case 2: ${\bf P}$ is supported on a family of Weibull distributions with fixed scale parameter $1$ and varying shape parameter $k$. This family of Weibull distributions is indexed by the shape parameter $k$ which takes value either $1$ or $2$ with equal probabilities, i.e., $k\sim\mathrm{Unif}\{1,2\}$. Let $Q$ be the Weibull distribution with shape parameter $k_{Q}$ ($k_{Q}$ equating either $1$ or $2$). Its theoretical WSD with respect to ${\bf P}$ is
    $\mathrm{SD}(Q;{\bf P})$ $=1-\sqrt{\int_{0}^{\infty}\Big{(}\mathbb{E}_{k\sim\mathrm{Unif}\{1,2\}}\frac{x-x^{k_{Q}/k}}{\mathcal{W}_{2}(k_{Q},k)}\Big{)}^{2}k_{Q}x^{k_{Q}-1}e^{-x^{k_{Q}}}\mathrm{d}x}$
    $=1-\sqrt{\int_{0}^{\infty}\Big{(}\frac{x-x^{k_{Q}/\overline{k}_{Q}}}{2\mathcal{W}_{2}(k_{Q},\overline{k}_{Q})}\Big{)}^{2}k_{Q}x^{k_{Q}-1}e^{-x^{k_{Q}}}\mathrm{d}x}$
    $=1-\frac{1}{2\mathcal{W}_{2}(k_{Q},\overline{k}_{Q})}\sqrt{\int_{0}^{\infty}\big{(}x-x^{k_{Q}/\overline{k}_{Q}}\big{)}^{2}k_{Q}x^{k_{Q}-1}e^{-x^{k_{Q}}}\mathrm{d}x}$
    $=1/2,$
    where the optimal map from $\mathrm{Weibull}(k_{Q})$ to $\mathrm{Weibull}(k)$ is
    $T_{k_{Q},k}(x)=F_{k}^{-1}\circ F_{k_{Q}}(x)=x^{k_{Q}/k},$
    using $\overline{k}_{Q}=3-k_{Q}$, the convention $/0=0$ 0, and where $\mathcal{W}_{2}(k_{Q},k)$ is derived by
    $\mathcal{W}_{2}(k_{Q},k)$ $=\sqrt{\int_{0}^{\infty}\big{(}x-x^{k_{Q}/k}\big{)}^{2}k_{Q}x^{k_{Q}-1}e^{-x^{k_{Q}}}\mathrm{d}x}.$
  • Case 3: ${\bf P}$ is supported on a family of isotropic bivariate Gaussian distributions with varying centers. The distribution of the Gaussian centers is supported on four points $\{\boldsymbol{\mu}_{1}=(1,0)^{\top},\boldsymbol{\mu}_{2}=(-1,0)^{\top},\boldsymbol{\mu}_{3}=(0,1)^{\top},\boldsymbol{\mu}_{4}=(0,-1)^{\top}\}$ with equal probabilities $1/4$. Let $Q$ be any one of $\mathcal{N}(\boldsymbol{\mu}_{k},\boldsymbol{I}_{2\times 2})$ for $k=1,\ldots,4$. The theoretical WSD is computed as, see Section 4.3,
    $\mathrm{SD}(Q;{\bf P})$ $=1-\Bigg{\|}\frac{1}{4}\sum_{k\neq q}\frac{\boldsymbol{\mu}_{q}-\boldsymbol{\mu}_{k}}{\|\boldsymbol{\mu}_{q}-\boldsymbol{\mu}_{k}\|}\Bigg{\|}=\frac{3-\sqrt{2}}{4}.$
  • Case 4: ${\bf P}$ is supported on a family of bivariate uniform distributions $\mathrm{Unif}\big{(}[0,c]^{2}\big{)}$ with $c\sim\mathrm{Unif}[1,2]$. Let $Q$ be $\mathrm{Unif}\big{(}[0,c_{q}]^{2}\big{)}$. Its theoretical WSD with respect to ${\bf P}$ is
    $\mathrm{SD}(Q;{\bf P})$ $=1-\sqrt{\int_{[0,c_{q}]^{2}}\bigg{\|}\mathbb{E}_{c\sim\mathrm{Unif}[1,2]}\ \frac{\boldsymbol{x}-(c/c_{q})\boldsymbol{x}}{\mathcal{W}_{2}(c_{q},c)}\bigg{\|}^{2}\frac{1}{c_{q}^{2}}\mathrm{d}\boldsymbol{x}}$
    $=1-\sqrt{\int_{[0,c_{q}]^{2}}\bigg{\|}\sqrt{3/2}\boldsymbol{x}\mathbb{E}_{c\sim\mathrm{Unif}[1,2]}\frac{1-(c/c_{q})}{|c_{q}-c|}\bigg{\|}^{2}\frac{1}{c_{q}^{2}}\mathrm{d}\boldsymbol{x}}$
    $=1-\sqrt{3(2-3/c_{q})^{2}/2\int_{[0,c_{q}]^{2}}\|\boldsymbol{x}\|^{2}\frac{1}{c_{q}^{2}}\mathrm{d}\boldsymbol{x}}$
    $=1-|2c_{q}-3|,$
    where the optimal map from $\mathrm{Unif}\big{(}[0,c_{q}]^{2}\big{)}$ to $\mathrm{Unif}\big{(}[0,c]^{2}\big{)}$ is the dilation
    $T_{c_{q},c}(\boldsymbol{x})=\frac{c}{c_{q}}\boldsymbol{x},$
    and $\mathcal{W}_{2}(c_{q},c)$ is computed as
    $\mathcal{W}_{2}(c_{q},c)=\sqrt{\int_{[0,c_{q}]^{2}}\|\boldsymbol{x}-(c/c_{q})\boldsymbol{x}\|^{2}(1/c_{q}^{2})\mathrm{d}\boldsymbol{x}}=\sqrt{2/3}|c_{q}-c|.$
For each of the above cases, we repeat independently the following experiment for $100$ times: generate the data array $\boldsymbol{X}$ via the two-stage sampling procedure in Section 6.2; then compute the empirical WSDs of the sampled distributions; finally, compare the theoretical WSD and the ensemble of $100$ empirical WSDs. We choose $m=1000,\ n=2000$. As shown in Figure 1, the empirical estimates are gathering tightly around the corresponding theoretical values.
Figure 1: The green solid line depicts the change of theoretical WSD along the parameter indexing ${\bf P}$. The black circles represent the distribution of empirical WSDs, with error bars indicating one standard deviation above and below the mean
8.2 Wasserstein spatial depth vs. conventional spatial depth
As discussed in Section 4, when $\mathbf{P}$ is supported on a location family, the WSD coincides with the spatial depth of the location parameter. We verify the equivalence between WSD and spatial depth in the four cases described below.
  • Case 1: $\mathbf{P}$ is supported on a set of $d=10$ dimensional Gaussian distributions with identity covariance matrix. The Gaussian centers are i.i.d. from $\mathrm{Unif}[-2,2]^{d}$.
  • Case 2: $\mathbf{P}$ is supported on a set of $d=10$ dimensional Gaussian distributions with a common covariance matrix. The common covariance matrix is chosen as $\Sigma_{i,j}=0.2^{|i-j|}$. The Gaussian centers are drawn in the same way as Case 1.
  • Case 3: The support of $\mathbf{P}$ is a set of uniform distributions on $d=10$ dimensional unite cubes with varying centers. The centers of the cubes are identically independently drawn from $\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{d\times d})$.
  • Case 4: $\mathbf{P}$ is supported on a set of univariate double exponential distributions with fixed rate equaling $1$ and varying locations. The locations are identically independently drawn from $\mathcal{N}(0,1)$.
The simulation procedure is as follows. First, $n=500$ distributions are drawn as described above in each case. Second, $m=500$ data points are randomly drawn for each sampled distribution. Third, the empirical WSD of each empirical distribution is computed according to (9), and the empirical spatial depths of the locations are computed as in (5). Finally, we check whether the empirical WSDs and spatial depths are approximately equal. As shown in Figure 2, there are nice equality relationships between the two depths.
Figure 2: The relationships between the WSD and conventional spatial depth in the four cases above.
8.3 Outlier detection
Like conventional statistical depth, WSD can be used to detect outlier distributions. We demonstrate its utility for outlier detection in two cases. In each case, we draw $n=500$ distributions from a population ${\bf P}$ and six outlier distributions which are relatively far away from the population. For each sampled distribution, we draw $m=500$ data points. All the distributions are on $\mathbb{R}^{d}$ with $d=10$.
  • Case 1: the population is a collection of Gaussian distributions with common identity covariance matrix and random centers, where the centers follow i.i.d. $\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$; the six outlier distributions are
    $\mathcal{N}((5,\ldots,5)^{\top},\boldsymbol{I}),\quad\mathcal{N}((5,\ldots,5)^{\top},\boldsymbol{\Sigma})\ \text{with}\ \Sigma_{i,j}=0.5^{|i-j|},$
    $\big{[}\mathrm{Gamma}(3,2)\big{]}^{d},\quad\big{[}\mathrm{Unif}(-6,6)\big{]}^{d},\quad\big{[}\mathrm{Beta}(0.1,0.1)\big{]}^{d},$
    $\mathrm{Multinomial}\big{(}2d,(0.25,0.25,0.15,0.15,0.15,0.01,0.01,0.01,0.01,0.01)\big{)}.$
    Here for a distribution $\mu$, $[\mu]^{d}$ is the distribution such that for $\boldsymbol{Z}=(\boldsymbol{Z}_{1},\ldots,\boldsymbol{Z}_{d})\sim[\mu]^{d}$ we have $\boldsymbol{Z}_{1},\ldots,\boldsymbol{Z}_{d}\sim_{\textit{i.i.d.}}\mu$.
  • Case 2: the population is a collection of uniform distributions $\big{[}\mathrm{Unif}(0,u)\big{]}^{d}$ with $u\sim\mathrm{Unif}(1,2)$; the outlier distributions are
    $\mathcal{N}\big{(}(3,\ldots,3)^{\top},\boldsymbol{I}\big{)},\quad\mathcal{N}\big{(}(-1,\ldots,-1)^{\top},\boldsymbol{\Sigma}\big{)}\ \text{with}\ \Sigma_{i,j}=0.5^{|i-j|},$
    $\big{[}\mathrm{Poisson}(3)\big{]}^{d},\quad\big{[}\mathrm{Binomial}(d,0.2)\big{]}^{d},\quad\big{[}\chi^{2}_{10}\big{]}^{d},$
[…]
For each case, we repeat similar experiments for $20$ times. The experimental procedure is as follows: draw the data array $\boldsymbol{X}$ according to Case 1 or Case 2; compute their empirical WSD according to (9); detect the outlier distributions whose empirical WSDs are smaller than the $1\%$ quantile of all the empirical WSDs. In each of these $20$ repetitions, all the outlier distributions can be detected for both Case 1 and Case 2. Figure 3 shows the result of one experiment, where the black and red dots represent distributions from ${\bf P}$ and the outlier distributions, respectively.
Figure 3: Left panel: the distributions are drawn according to Case 1. Right panel: the distributions are drawn according to Case 2. The red dots represent the outlier distributions.
9 Application
Nowadays, climate change is a major concern across the society. Considerable amount of information can be extracted from longitudinal series of daily temperatures. We apply the notion of WSD to explore a dataset recording European daily temperatures during the past two centuries.
The data is collected from the public database “European Climate Assessment and Dataset”111 https://www.ecad.eu/dailydata/index.php.. It contains the daily average temperatures collected at $40$ meteorological stations located across Europe, including Austria, Croatia, Czech Republic, Denmark, Finland, Germany, Sweden, and United Kingdom, from year 1874 to 2023. These $40$ meteorological stations cover a broad range of Europe and are representative of the region. The goal is to explore the temperature change over the years.
We consider monthly temperatures obtained by averaging daily temperatures per month. To each weather station, we obtain a $12$ monthly-average temperature curve, represented by a vector in $\mathbb{R}^{12}$. Hence, the monthly temperatures of each year correspond to one distribution on $\mathbb{R}^{12}$. For a particular year, the $12$ monthly temperatures (forming a vector in $\mathbb{R}^{12}$) collected at each station act as a sample point drawn from this distribution. Finally, we gather $150$ distributions (from year 1874 to year 2023) with each distribution associated to $40$ sample points (for the $40$ meteorological stations), and where each sample point is a $12$-dimensional real vector. Contrary to other work, we do not consider the annual evolution of the temperatures for a particular place but rather analyze the different temperature curves at all locations at the same time. We aim at understanding weather change at a global scale by considering the $40$ different locations as representatives of the European climate.
Within this framework, we compute the empirical WSDs of these $150$ distributions as in (9). Several outlier years are identified based on their excessively small WSDs. As we discuss next, these identified “abnormal years” are consistent with historical records, which further validate the practical utility of Wasserstein spatial depth.
For the reproducibility of our research, the code for data analysis is publicly available at https://github.com/YishaYao/Wasserstein-Spatial-Depth/tree/main.
Figure 4: The green dots represent regular/representative/central distributions, while the red dots correspond to the distributions near the outskirt and “far” from the center
The values of the $150$ empirical WSDs are shown in Figure 4. The lowest $5\%$ values, which we consider as outliers, are colored red, and the corresponding years are also marked. Based on empirical WSDs, the temperatures at years 1879, 1929, 1940, 1942, 1947, 1956, 1963, and 2018 are more “exotic” or near outskirt. After searching among historical documentations, we indeed fond evidences to support this discovery. Year 1879 was an extremely cold year, featured with a unusually snowy winter (November and December); the first two months of 1929 were recorded as one of the coldest winters in Europe during the past century with temperature reaching down to -30°C in central Europe. Both year 1940 and year 1942 were marked by severe winters with dramatic ice storms, and year 1942 had a cool summer; The weather in year 1947 was unusually cold in winter and record-breaking hot in summer; Europe experienced severe cold waves in both winters of 1956 and 1963; The well-known 2018 European drought and heat wave led to record-breaking temperatures and wildfires in many parts of Europe.
To get a better view on how these years’ temperatures differ from other regular years’, we compare the four most “exotic” years with the most regular years. We pick the two years with the largest WSDs as our “regular years”, year 1935 and year 1960. In each plot of Figure 5, the bundle of green curves represents the temperature trends of the $40$ locations in the regular years (1935 and 1960), while the bundle of red curves corresponds to one particular outlier year. The green bundle and red bundle do exhibit clear visual differences in temperature trends over the months.
(a) outlier year 1929
(b) outlier year 1940
(c) outlier year 1956
(d) outlier year 2018
Figure 5: Comparisons between the most regular years 1935 and 1960 (the two years with the largest Wasserstein spatial depths) and four outlier years. In each plot, the bundle of green curves represents the temperature trends in years 1935 and 1960 at 40 locations in Europe (totally $80$ green curves); the bundle of red curves represents the temperature trends in an outlier year at the same 40 locations (totally $40$ red curves).
10 Further directions and future work
In this work, we propose a new notion of depth on the Wasserstein space. We demonstrate that it preserves critical properties of conventional statistical depths. Additionally, it has a straightforward empirical counterpart that can be easily computed from sample data and is asymptotically consistent. Numerical simulations and real data analysis further support its practical utility.
Note that we have defined the new notion of Wasserstein spatial depth $\mathrm{SD}(Q;{\bf P})$ for absolutely continuous distributions $Q$ and where ${\bf P}$ can be arbitrary. This is because our approach exploits the definition of the geodesics in the Wasserstein space (see Section 3.3).
When $Q$ is not absolutely continuous, the geodesic between $Q$ and another distribution $P$ might be not unique. In this case, the set of geodesics is given by the laws of the random vectors $(1-t)\mathbf{X}+t\mathbf{Y}$ where the law of the random vector $(\mathbf{X},\mathbf{Y})$, namely $\pi_{P,Q}$, is an optimal transport plan, as in (1) with $p=2$. Hence, uniqueness of the geodesics is equivalent to uniqueness of the transport plans.
Thus, if with ${\bf P}$-probability one $P\sim{\bf P}$ is absolutely continuous, even if $Q$ is not absolutely continuous, the geodesics are unique and, following the route of Section 3.3, we can still define a notion of depth as follows:
[…]
where $\pi_{Q,P,P^{\prime}}(\mathbf{x},\mathbf{y},\mathbf{y}^{\prime})$ is the distribution of a vector $(\mathbf{X},\mathbf{Y},\mathbf{Y}^{\prime})$ with $(\mathbf{X},\mathbf{Y})\sim\pi_{Q,P}$, $(\mathbf{X},\mathbf{Y}^{\prime})\sim\pi_{Q,P^{\prime}}$ and $\mathbf{Y}$ and $\mathbf{Y}^{\prime}$ are independent given $\mathbf{X}$. Here $\pi_{Q,P}$ (resp. $\pi_{Q,P^{\prime}}$) is the unique optimal transport plan from $Q$ to $P$ (resp. $P^{\prime}$). This provides a definition of WSD for any distribution $Q$ when ${\bf P}$ samples a.s. absolutely continuous distributions. It can be seen, similarly as the proof of Theorem 5.1, that $\mathrm{SD}^{\text{discr}}(Q;{\bf P})$ would be $[0,1]$-valued (and the quantity in the square root being non-negative).
We leave for future exploration the practical utility of this complementary WSD, along with the task of establishing analogous favorable mathematical properties as those demonstrated in this paper. Note that the depth in (10) coincides with $\mathrm{SD}(Q;{\bf P})$ in the special case where both $Q$ and (a.s.) the samples from P are absolutely continuous. This can be seen from the arguments leading to (A.3).
Finally, for computational reasons, the statistics and machine learning community has also focused on regularized optimal transport [15, 35]. It is an interesting prospect as well to extend the Wasserstein spatial depth to regularized optimal transport.
Appendix A Proof of Theorem 5.1
A.1 Values in $[0,1]$
Here we prove that $\mathrm{SD}(Q;{\bf P})\in[0,1]$ for all $Q\in\mathcal{P}^{a.c}_{2}(\mathbb{R}^{d})$, which is probably the easiest statement to prove. To prove the upper bound we realize that
$\left(\int{\left\|\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{\mathbf{x}-T_{Q,P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}\right]\right\|^{2}}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}}\geq 0,$
so that
$\mathrm{SD}(Q;{\bf P})=1-\left(\int{\left\|\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{\mathbf{x}-T_{Q,P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}\right]\right\|^{2}}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}}\leq 1.$
To prove the lower bound we observe that
$\left(\int{\left\|\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{\mathbf{x}-T_{Q,P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}\right]\right\|^{2}}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}}$
$=\sup_{\|G\|_{L^{2}(Q)}\leq 1}\left(\int{\left\langle\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{\mathbf{x}-T_{Q,P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}\right],G(\mathbf{x})\right\rangle}\mathrm{d}Q(\mathbf{x})\right)$
$=\sup_{\|G\|_{L^{2}(Q)}\leq 1}\left(\int{\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{\left\langle\mathbf{x}-T_{Q,P}(\mathbf{x}),G(\mathbf{x})\right\rangle}{\mathcal{W}_{2}(P,Q)}\right]}\mathrm{d}Q(\mathbf{x})\right)$
$\leq\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{\sup_{\|G\|_{L^{2}(Q)}\leq 1}\int{\left\langle\mathbf{x}-T_{Q,P}(\mathbf{x}),G(\mathbf{x})\right\rangle}\mathrm{d}Q(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}\right]$
$=\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{\|I-T_{Q,P}\|_{L^{2}(Q)}}{\mathcal{W}_{2}(P,Q)}\right]=\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{\mathcal{W}_{2}(P,Q)}{\mathcal{W}_{2}(P,Q)}\right]=1,$
so that
$\mathrm{SD}(Q;{\bf P})=1-\left(\int{\left\|\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{\mathbf{x}-T_{Q,P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}\right]\right\|^{2}}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}}\geq 0.$
A.2 Transformation invariance
Theorem 1.2 in [23] describes the group of isometries of $(\mathcal{P}_{2}(\mathbb{R}^{d}),\mathcal{W}_{2})$ for $d>2$. Any isometry $F$, can be written as the composition of $\Phi(\varphi)$ and a trivial isometry. Recall that $\Phi(\varphi):P\mapsto\Phi(\varphi)(P)$ where $\varphi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a linear isometry and $\Phi(\varphi)(P)$ is the law of the random variable
$\varphi(\mathbf{X}-\mathbb{E}[\mathbf{X}])+\mathbb{E}[\mathbf{X}],\quad{\rm for}\ \mathbf{X}\sim P.$
Therefore, it is enough to show that the WSD is invariant with respect to trivial isometries and isometries of type $\Phi(\varphi)$ for some linear isometry $\varphi:\mathbb{R}^{d}\to\mathbb{R}^{d}$.
Invariance under trivial isometries.
Let $\mathbf{A}$ be a $d\times d$ orthogonal matrix and ${\bf b}\in\mathbb{R}^{d}$. We write
$f_{\mathbf{A},{\bf b}}(\mathbf{x})=\mathbf{A}\mathbf{x}+{\bf b}.$
The mapping
$S_{P}=f_{\mathbf{A},{\bf b}}\circ T_{Q,P}\circ(f_{\mathbf{A},{\bf b}})^{-1}:\mathbf{x}\mapsto\mathbf{A}\,T_{Q,P}(\mathbf{A}^{T}(\mathbf{x}-{\bf b}))+{\bf b}$
is the a.s. defined gradient of a convex function and (by construction) pushes $(f_{\mathbf{A},{\bf b}})_{\#}Q$ forward to $(f_{\mathbf{A},{\bf b}})_{\#}P$. Therefore, $S_{P}$ is the optimal transport map from $(f_{\mathbf{A},{\bf b}})_{\#}Q$ forward to $(f_{\mathbf{A},{\bf b}})_{\#}P$ (cf. [28]). Hence, the following holds for the induced isometry $F:P\mapsto F(P)=(f_{\mathbf{A},{\bf b}})_{\#}P$:
$\mathrm{SD}(F(Q);F_{\#}{\bf P})$ $=1-\left(\int{\left\|\mathbb{E}_{P}\left[\frac{\mathbf{x}-S_{P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}\right]\right\|^{2}}\mathrm{d}((f_{\mathbf{A},{\bf b}})_{\#}Q)(\mathbf{x})\right)^{\frac{1}{2}}$
$=1-\left(\int{\left\|\mathbb{E}_{P}\left[\frac{f_{\mathbf{A},{\bf b}}(\mathbf{x})-f_{\mathbf{A},{\bf b}}\circ T_{Q,P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}\right]\right\|^{2}}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}}$
$=1-\left(\int{\left\|\mathbb{E}_{P}\left[\frac{\mathbf{A}(\mathbf{x}-T_{Q,P}(\mathbf{x}))}{\mathcal{W}_{2}(P,Q)}\right]\right\|^{2}}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}}$
$=1-\left(\int{\left\|\mathbf{A}\,\mathbb{E}_{P}\left[\frac{\mathbf{x}-T_{Q,P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}\right]\right\|^{2}}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}}$
$=1-\left(\int{\left\|\mathbb{E}_{P}\left[\frac{\mathbf{x}-T_{Q,P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}\right]\right\|^{2}}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}}=\mathrm{SD}(Q;{\bf P}).$
This proves the invariance under trivial isometries. ∎
Invariance under isometries of type $\Phi(\varphi)$.
Let $\varphi$ be a linear isometry. Then the mapping $S_{P}$ solving
$S_{P}(\varphi(\mathbf{x}-\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}])+\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}]))=\varphi(T_{Q,P}(\mathbf{x})-\mathbb{E}_{\mathbf{Y}\sim P}[\mathbf{Y}])+\mathbb{E}_{\mathbf{Y}\sim P}[\mathbf{Y}])$
is, as in the previous case, the optimal transport map from $\Phi(\varphi)(Q)$ to $\Phi(\varphi)(P)$. Then it holds that
$\mathrm{SD}(\Phi(\varphi)(Q);(\Phi(\varphi))_{\#}{\bf P})$ $=1-\left(\int{\left\|\mathbb{E}_{P\sim{\bf P}}\left[\frac{\mathbf{x}-S_{P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}\right]\right\|^{2}}\mathrm{d}\Phi(\varphi)(Q))(\mathbf{x})\right)^{\frac{1}{2}}$
$=1-\bigg{(}\int\bigg{\|}\mathbb{E}_{P\sim{\bf P}}\bigg{[}\frac{\varphi(\mathbf{x}-\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}])+\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}]))}{\mathcal{W}_{2}(P,Q)}$
$\qquad\qquad-\frac{\varphi(T_{Q,P}(\mathbf{x})-\mathbb{E}_{\mathbf{Y}\sim P}[\mathbf{Y}])+\mathbb{E}_{\mathbf{Y}\sim P}[\mathbf{Y}]}{\mathcal{W}_{2}(P,Q)}\bigg{]}\bigg{\|}^{2}\mathrm{d}Q(\mathbf{x})\bigg{)}^{\frac{1}{2}}.$
As $\varphi$ is linear, we get the equality
$\mathrm{SD}(\Phi(\varphi)(Q);(\Phi(\varphi))_{\#}{\bf P})=$ $1-\bigg{(}\int\bigg{\|}\mathbb{E}_{P\sim{\bf P}}\bigg{[}\frac{\varphi(\mathbf{x}-T_{Q,P}(\mathbf{x})-\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}]+\mathbb{E}_{\mathbf{Y}\sim P}[\mathbf{Y}])}{\mathcal{W}_{2}(P,Q)}$
$\qquad\qquad\qquad\qquad+\frac{\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}]-\mathbb{E}_{\mathbf{Y}\sim P}[\mathbf{Y}]}{\mathcal{W}_{2}(P,Q)}\bigg{]}\bigg{\|}^{2}\mathrm{d}Q(\mathbf{x})\bigg{)}^{\frac{1}{2}}$
$1-\bigg{(}\int\bigg{\|}\varphi\left(\mathbb{E}_{P\sim{\bf P}}\bigg{[}\frac{\mathbf{x}-T_{Q,P}(\mathbf{x})-\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}]+\mathbb{E}_{\mathbf{Y}\sim P}[\mathbf{Y}])}{\mathcal{W}_{2}(P,Q)}\bigg{]}\right)$
$\qquad\qquad\qquad\qquad+\mathbb{E}_{P\sim{\bf P}}\bigg{[}\frac{\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}]-\mathbb{E}_{\mathbf{Y}\sim P}[\mathbf{Y}]}{\mathcal{W}_{2}(P,Q)}\bigg{]}\bigg{\|}^{2}\mathrm{d}Q(\mathbf{x})\bigg{)}^{\frac{1}{2}}.$
Develop the squares and use the fact that $\varphi$ is an isometry to obtain
$\mathrm{SD}(\Phi(\varphi)(Q);(\Phi(\varphi))_{\#}{\bf P})$
$=$ $1-\bigg{(}\int\bigg{\{}\bigg{\|}\mathbb{E}_{P\sim{\bf P}}\bigg{[}\frac{\mathbf{x}-T_{Q,P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}\bigg{]}\bigg{\|}^{2}+2\bigg{\|}\mathbb{E}_{P\sim{\bf P}}\bigg{[}\frac{\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}]-\mathbb{E}_{\mathbf{Y}\sim P}[\mathbf{Y}]}{\mathcal{W}_{2}(P,Q)}\bigg{]}\bigg{\|}^{2}$
$+2\bigg{\langle}\mathbb{E}_{P\sim{\bf P}}\bigg{[}\frac{\mathbf{x}-T_{Q,P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}\bigg{]},\mathbb{E}_{P\sim{\bf P}}\bigg{[}\frac{\mathbb{E}_{\mathbf{Y}\sim P}[\mathbf{Y}]-\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}]}{\mathcal{W}_{2}(P,Q)}\bigg{]}\bigg{\rangle}$
$-2\bigg{\langle}\varphi\left(\mathbb{E}_{P\sim{\bf P}}\bigg{[}\frac{\mathbf{x}-T_{Q,P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}\bigg{]}\right),\mathbb{E}_{P\sim{\bf P}}\bigg{[}\frac{\mathbb{E}_{\mathbf{Y}\sim P}[\mathbf{Y}]-\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}]}{\mathcal{W}_{2}(P,Q)}\bigg{]}\bigg{\rangle}$
$-2\bigg{\langle}\varphi\left(\mathbb{E}_{P\sim{\bf P}}\bigg{[}\frac{\mathbb{E}_{\mathbf{Y}\sim P}[\mathbf{Y}]-\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}]}{\mathcal{W}_{2}(P,Q)}\bigg{]}\right),\mathbb{E}_{P\sim{\bf P}}\bigg{[}\frac{\mathbb{E}_{\mathbf{Y}\sim P}[\mathbf{Y}]-\mathbb{E}_{\mathbf{X}\sim Q}[\mathbf{X}]}{\mathcal{W}_{2}(P,Q)}\bigg{]}\bigg{\rangle}\bigg{\}}\mathrm{d}Q(\mathbf{x})\bigg{)}^{\frac{1}{2}}.$
The second term of the sum cancels with the third and the fourth with the last one as a consequence of Fubini’s theorem, the linearity of $\varphi$ and the fact that $(T_{Q,P})_{\#}Q=P$. Therefore, the result follows. ∎
A.3 Vanishing at infinity
The goal of this section is to prove that $\mathrm{SD}(Q_{n};{\bf P})\to 0$ as $\mathcal{W}_{2}(Q_{n},P)\to\infty$ for one $P\in\mathcal{P}_{2}(\mathbb{R}^{d})$.
Remark A.1.
Note that $\mathcal{W}_{2}(Q_{n},P)\to\infty$ implies that for any other $P^{\prime}\in\mathcal{P}_{2}(P)$,
$\mathcal{W}_{2}(Q_{n},P^{\prime})\geq\mathcal{W}_{2}(Q_{n},P)-\mathcal{W}_{2}(P^{\prime},P)\to+\infty.$
Moreover, for any compact set $K$,
$\inf_{P\in K}\mathcal{W}_{2}(Q_{n},P)\to+\infty.$
Let $\{Q_{n}\}_{n\in\mathbb{N}}\subset\mathcal{P}^{a.c}_{2}(\mathbb{R}^{d})$ be such that $\mathcal{W}_{2}(Q_{n},P)\to\infty$ for all $P\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Recall that
$\mathrm{SD}(Q_{n};{\bf P}):=1-\left(\int{\left\|\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{\mathbf{x}-T_{Q_{n},P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q_{n})}\right]\right\|^{2}}\mathrm{d}Q_{n}(\mathbf{x})\right)^{\frac{1}{2}}$
with the convention $\frac{\mathbf{x}-T_{Q_{n},P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q_{n})}=0$ if ${\mathcal{W}_{2}(P,Q_{n})}=0$. First we want to get rid of this last pathological case. Let
$A_{n}:=\int{\left\|\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{\mathbf{x}-T_{Q_{n},P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q_{n})}\right]\right\|^{2}}\mathrm{d}Q_{n}(\mathbf{x}).$ (11)
Let $E_{n}=\{Q_{n}\}$. Note that when $P\in E_{n}$, a 00 appears in the expression of $A_{n}$ (recall the convention $/0=0$ 0). For each $n$, we modify ${\bf P}={\bf P}_{1}+{\bf P}_{2}$, where ${\bf P}_{1}$ is a measure on $\mathcal{P}_{2}(\mathbb{R}^{d})\backslash E_{n}$ and ${\bf P}_{2}$ is a measure on $E_{n}$, by ${\bf P}^{\prime}={\bf P}_{1}+{\bf\widetilde{P}}_{2}$, where ${\bf\widetilde{P}}_{2}$ is an arbitrary measure on $\mathcal{P}_{2}(\mathbb{R}^{d})\backslash E_{n}$ such that ${\bf\widetilde{P}}_{2}(\mathcal{P}_{2}(\mathbb{R}^{d}))={\bf P}(E_{n})$. Note that ${\bf P}^{\prime}$ is also a probability measure.
Since the measure ${\bf P}$ is tight and $Q_{n}$ diverges, it is clear that ${\bf P}(E_{n})\to 0$ as $n\to\infty$. Moreover,
$\left|\left(\int{\left\|\mathbb{E}_{P\sim\mathbf{P}^{\prime}}\left[\frac{\mathbf{x}-T_{Q_{n},P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q_{n})}\right]\right\|^{2}}\mathrm{d}Q_{n}(\mathbf{x})\right)^{1/2}-(A_{n})^{1/2}\right|$
$\leq{\bf P}(E_{n})\left(\int{\left\|\mathbb{E}_{P\sim\frac{{\bf\widetilde{P}}_{2}}{{\bf P}(E_{n})}}\left[\frac{\mathbf{x}-T_{Q_{n},P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q_{n})}\right]\right\|^{2}}\mathrm{d}Q_{n}(\mathbf{x})\right)^{1/2}.$
Since the spatial depth lies in $[0,1]$, we can upper bound this quantity by ${\bf P}(E_{n})$ and obtain that the limit of $\mathrm{SD}(Q_{n};{\bf P})$ is that of $\mathrm{SD}(Q_{n};{\bf P}^{\prime})$. Therefore, we can feel free to assume that ${\mathcal{W}_{2}(P,Q_{n})}=0$ does not happen for $n$ big enough and for $P\sim{\bf P}$.
We prove that $A_{n}\to 1$, where $A_{n}$ is defined in (11). To do so, let $P^{\prime}$ be an independent copy of $P$, so that
$A_{n}$ $=\int{\left\langle\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{\mathbf{x}-T_{Q_{n},P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q_{n})}\right],\mathbb{E}_{P^{\prime}\sim\mathbf{P}}\left[\frac{\mathbf{x}-T_{Q_{n},P^{\prime}}(\mathbf{x})}{\mathcal{W}_{2}(P^{\prime},Q_{n})}\right]\right\rangle}\mathrm{d}Q_{n}(\mathbf{x})$
$=\int{\mathbb{E}_{P,P^{\prime}\sim\mathbf{P}}\left[\frac{\left\langle\mathbf{x}-T_{Q_{n},P}(\mathbf{x}),\mathbf{x}-T_{Q_{n},P^{\prime}}(\mathbf{x})\right\rangle}{\mathcal{W}_{2}(P,Q_{n})\mathcal{W}_{2}(P^{\prime},Q_{n})}\right]}\mathrm{d}Q_{n}(\mathbf{x}).$ (12)
In order to reduce the size of the formulas we call $B_{P,n}(\mathbf{x})=\mathbf{x}-T_{Q_{n},P}(\mathbf{x})$ and $B_{P^{\prime},n}(\mathbf{x})=\mathbf{x}-T_{Q_{n},P^{\prime}}(\mathbf{x})$. Then
$A_{n}=\int{\mathbb{E}_{P,P^{\prime}\sim\mathbf{P}}\left[\frac{\left\langle B_{P,n}(\mathbf{x}),B_{P^{\prime},n}(\mathbf{x})\right\rangle}{\|B_{P,n}\|_{L^{2}(Q_{n})}\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\right]}\mathrm{d}Q_{n}(\mathbf{x}),$
and, via Fubini’s theorem,
$A_{n}=\mathbb{E}_{P,P^{\prime}\sim\mathbf{P}}\left[\frac{\left\langle B_{P,n},B_{P^{\prime},n}\right\rangle_{L^{2}(Q_{n})}}{\|B_{P,n}\|_{L^{2}(Q_{n})}\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\right].$
Since
$C_{n}(P,P^{\prime}):=\frac{\left\langle B_{P,n},B_{P^{\prime},n}\right\rangle_{L^{2}(Q_{n})}}{\|B_{P,n}\|_{L^{2}(Q_{n})}\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\leq\frac{\|B_{P,n}\|_{L^{2}(Q_{n})}\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}{\|B_{P,n}\|_{L^{2}(Q_{n})}\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}=1,$
the dominated convergence theorem can be applied and we only need to show that
$C_{n}(P,P^{\prime})\longrightarrow 1,\quad\text{for }{\bf P}\otimes{\bf P}-{\rm a.e.}\ (P,P^{\prime}).$ (13)
We decompose $C_{n}(P,P^{\prime})$ in two terms: $C_{n}(P,P^{\prime})=C_{n,1}(P,P^{\prime})+C_{n,2}(P,P^{\prime})$ with
$C_{n,1}(P,P^{\prime})=\frac{\|B_{P,n}\|_{L^{2}(Q_{n})}^{2}}{\|B_{P,n}\|_{L^{2}(Q_{n})}^{2}}=1,$
and
$C_{n,2}(P,P^{\prime})=\left\langle\frac{B_{P,n}}{\|B_{P,n}\|_{L^{2}(Q_{n})}},\frac{B_{P^{\prime},n}}{\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}-\frac{B_{P,n}}{\|B_{P,n}\|_{L^{2}(Q_{n})}}\right\rangle_{L^{2}(Q_{n})}.$
The goal, of course, is to show that $C_{n,2}(P,P^{\prime})\to 0$, for ${\bf P}\otimes{\bf P}$-a.e. $(P,P^{\prime})$. Since
$C_{n,2}(P,P^{\prime})$
$=\left\langle\frac{B_{P,n}}{\|B_{P,n}\|_{L^{2}(Q_{n})}},\frac{B_{P^{\prime},n}-B_{P,n}}{\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}+B_{P,n}\left(\frac{1}{\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}-\frac{1}{\|B_{P,n}\|_{L^{2}(Q_{n})}}\right)\right\rangle_{L^{2}(Q_{n})}$
$=\left\langle\frac{B_{P,n}}{\|B_{P,n}\|_{L^{2}(Q_{n})}},\frac{T_{Q_{n},P}-T_{Q_{n},P^{\prime}}}{\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}+B_{P,n}\left(\frac{1}{\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}-\frac{1}{\|B_{P,n}\|_{L^{2}(Q_{n})}}\right)\right\rangle_{L^{2}(Q_{n})},$
we can upper bound $|C_{n,2}(P,P^{\prime})|$ by
$\frac{\|B_{P,n}\|_{L^{2}(Q_{n})}}{\|B_{P,n}\|_{L^{2}(Q_{n})}}\frac{\|T_{Q_{n},P}\|_{L^{2}(Q_{n})}+\|T_{Q_{n},P^{\prime}}\|_{L^{2}(Q_{n})}}{\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\\ +\left|\left\langle\frac{B_{P,n}}{\|B_{P,n}\|_{L^{2}(Q_{n})}},B_{P,n}\left(\frac{1}{\|B_{P,n}\|_{L^{2}(Q_{n})}}-\frac{1}{\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\right)\right\rangle_{L^{2}(Q_{n})}\right|.$ (14)
The first term of (14) tends to 00 for ${\bf P}\otimes{\bf P}$-a.e. $(P,P^{\prime})$. Indeed, using Cauchy-Schwarz inequality, $(a+b)^{2}\leq 2(a^{2}+b^{2})$ and the equality $\|T_{Q_{n},P}\|_{L^{2}(Q_{n})}^{2}=\int\|\mathbf{x}\|^{2}\mathrm{d}P(\mathbf{x})$, the following upper bound
$\frac{\int\|\mathbf{x}\|^{2}\mathrm{d}P(\mathbf{x})+\int\|\mathbf{x}\|^{2}\mathrm{d}P^{\prime}(\mathbf{x})}{\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}$ (15)
holds for the first term of (14). The latter clearly tends to 00 since $\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}=\mathcal{W}_{2}(Q_{n},P^{\prime})$.
To show that the second term of (14) also tends to 00 we use the bound
$\left|\left\langle\frac{B_{P,n}}{\|B_{P,n}\|_{L^{2}(Q_{n})}},B_{P,n}\left(\frac{1}{\|B_{P,n}\|_{L^{2}(Q_{n})}}-\frac{1}{\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\right)\right\rangle_{L^{2}(Q_{n})}\right|\\ \leq\left|\|B_{P,n}\|\left(\frac{\|B_{P,n}\|_{L^{2}(Q_{n})}-\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}|}{\|B_{P,n}\|_{L^{2}(Q_{n})}\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\right)\right|$
followed by the triangle inequality
$\left|\|B_{P,n}\|\left(\frac{\|B_{P,n}\|_{L^{2}(Q_{n})}-\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}|}{\|B_{P,n}\|_{L^{2}(Q_{n})}\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\right)\right|$ $=\left(\frac{\|B_{P,n}\|_{L^{2}(Q_{n})}-\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}|}{\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\right)$
$\leq\left(\frac{\|B_{P,n}-B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}{\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\right)$
$=\left(\frac{\|T_{Q_{n},P}-T_{Q_{n},P^{\prime}}\|_{L^{2}(Q_{n})}}{\|B_{P^{\prime},n}\|_{L^{2}(Q_{n})}}\right).$
The latter can be upper bounded by (15), so that the second term of (14) also tends to 00 for ${\bf P}\otimes{\bf P}$-a.e. $(P,P^{\prime})$. This implies $C_{n,2}(P,P^{\prime})$ tends to 00 for ${\bf P}\otimes{\bf P}$-a.e. $(P,P^{\prime})$. Hence, (13) holds.
Appendix B Proof of Theorem 5.2
Consider $P\in\{P_{1},\ldots,P_{n}\}$ and $Q\in\mathcal{P}_{2}^{a.c.}(\mathbb{R}^{d})$ such that $Q\not\in\{P_{1},\ldots,P_{n}\}$. We recall from [42] that
$\mathcal{W}_{2}^{2}(P,Q)=\inf_{\pi\in\Pi(P,Q)}\frac{1}{2}\int\|\mathbf{x}-\mathbf{y}\|^{2}\mathrm{d}\pi(\mathbf{x},\mathbf{y})$ (16)
admits a dual formulation
$\mathcal{W}_{2}^{2}(P,Q)=\sup_{(f,g)\in\Phi}\left\{\int f(\mathbf{x})\,\mathrm{d}Q(\mathbf{x})+\int g(\mathbf{y})\,\mathrm{d}P(\mathbf{y})\right\},$ (17)
where $\Phi=\{(f,g)\in\mathcal{C}(\mathbb{R}^{d})\times\mathcal{C}(\mathbb{R}^{d}):\ f(\mathbf{x})+g(\mathbf{y})\leq\frac{1}{2}\|\mathbf{x}-\mathbf{y}\|^{2}\}$. Here $\mathcal{C}(\mathbb{R}^{d})$ is the set of continuous functions on $\mathbb{R}^{d}$. We denote as $(f_{Q,P},f_{P,Q})$ the solutions of (17). It is well-known that $\nabla f_{Q,P}(\mathbf{x})=\mathbf{x}-T_{Q,P}(\mathbf{x})$. The indicator function of a set $A$ is denoted as ${\bf 1}_{A}$. Now we argue by contradiction. We assume first that there exists
$Q\in\mathcal{P}_{2}^{a.c}(\mathbb{R}^{d})\cap\operatorname*{arg\,min}_{P}\mathbb{E}_{P\sim\mathbf{P}}[\mathcal{W}_{2}(P,Q)]$
with $Q\not\in\{P_{1},\ldots,P_{n}\}$ and we assume that the set $\mathcal{K}$ of all $\mathbf{x}$ such that
$s(\mathbf{x}):=\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{\mathbf{x}-T_{Q,P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}\right]\neq 0$
has positive measure $Q(\mathcal{K})>0$. As $T_{Q,P}$ is the gradient of a l.s.c. convex function, it is continuous $Q$-a.e., so that $s$ is also continuous $Q$-a.e. Therefore, there exists a compact convex set with non-empty interior $U$ such that $U\subset\mathcal{K}$. Consider the signed measure $h$ such that $\frac{dh}{dQ}=-{\bf 1}_{U}\left(f_{Q,P}-\int_{U}f_{Q,P}(\mathbf{z})\mathrm{d}Q(\mathbf{z})\right)$. Note that $h(\mathbb{R}^{d})=0$ and $Q+th$ is a probability measure with finite second order moment for all $t$ in a neighborhood of zero. Since $(\cdot)^{1/2}$ is concave,
$\mathcal{W}_{2}(P,Q+th)\leq\mathcal{W}_{2}(P,Q)+\frac{\mathcal{W}_{2}^{2}(P,Q+th)-\mathcal{W}_{2}^{2}(P,Q)}{2\mathcal{W}_{2}(P,Q)}.$
Using the dual formulation (17) we obtain for $t$ in a neighborhood of zero,
$\frac{\mathcal{W}_{2}(P,Q+th)-\mathcal{W}_{2}(P,Q)}{t}\leq\frac{\int f_{Q+th,P}(\mathbf{x})\,\mathrm{d}h(\mathbf{x})}{2\mathcal{W}_{2}(P,Q)}.$
Since $h(\mathbb{R}^{d})=0$, we have for $t$ in a neighborhood of zero
$\frac{\mathcal{W}_{2}(P,Q+th)-\mathcal{W}_{2}(P,Q)}{t}\\ \leq-\frac{\int_{U}(f_{Q,P}(\mathbf{x})-\int_{U}f_{Q,P}(\mathbf{z})\mathrm{d}Q(\mathbf{z}))(f_{Q+th,P}(\mathbf{x})-\int_{U}f_{Q+th,P}(\mathbf{z})\mathrm{d}Q(\mathbf{z}))\mathrm{d}Q(\mathbf{x})}{2\mathcal{W}_{2}(P,Q)}.$
Set
$M(P):=\frac{1}{2}\frac{\int_{U}(f_{Q,P}(\mathbf{x})-\int_{U}f_{Q,P}(\mathbf{z})\mathrm{d}Q(\mathbf{z}))^{2}\mathrm{d}Q(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}$
and the norm
$\|\phi\|_{U}:=\left(\int_{U}\left(\phi(\mathbf{x})-\int_{U}\phi(\mathbf{z})\mathrm{d}Q(\mathbf{z})\right)^{2}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}}.$
Then
$\frac{\mathcal{W}_{2}(P,Q+th)-\mathcal{W}_{2}(P,Q)}{t}\leq-M(P)+\frac{\|f_{Q,P}\|_{U}\|f_{Q,P}-f_{Q+th,P}\|_{U}}{2\mathcal{W}_{2}(P,Q)}.$
Since $s(\mathbf{x})\neq 0$ for $\mathbf{x}\in U$, the function $U\ni\mathbf{x}\mapsto\mathbb{E}[f_{Q,P}(\mathbf{x})]$ is non constant, which implies that
$\mathbb{E}_{P\sim{\bf P}}[M(P)]:=\frac{1}{2}\mathbb{E}_{P\sim{\bf P}}\left[\frac{\int_{U}(f_{Q,P}(\mathbf{x})-\int_{U}f_{Q,P}(\mathbf{z})\mathrm{d}Q(\mathbf{z}))^{2}\mathrm{d}Q(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}\right]>0.$
The theorem follows upon showing that
$\mathbb{E}_{P\sim{\bf P}}\left[\frac{\|f_{Q,P}\|_{U}\|f_{Q,P}-f_{Q+th,P}\|_{U}}{\mathcal{W}_{2}(P,Q)}\right]\to 0\quad{\rm as}\ t\to 0,$ (18)
which is a trivial consequence of the main result of [36] and the assumption $Q\not\in\{P_{1},\ldots,P_{n}\}$.
Appendix C Proofs of Section 5.3
Proof of Theorem 5.4.
As $\mathcal{W}_{2}(Q_{n},Q)\to 0$ and $\mathbf{P}\in\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ is atomless there exists an open Wasserstein ball
$\mathbb{B}_{\mathcal{W}_{2}}(Q,\beta)=\{P\in\mathcal{P}_{2}(\mathbb{R}^{d}):\ \mathcal{W}_{2}(P,Q)<\beta\}$
with ${\bf P}(\mathbb{B}_{\mathcal{W}_{2}}(Q,\beta))\leq\epsilon/2$. Since ${\bf P}\in\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ is tight, there exists a compact set $K\subset\mathcal{P}_{2}(\mathbb{R}^{d})$ such that ${\bf P}(\mathcal{P}_{2}(\mathbb{R}^{d})\setminus K)\leq\epsilon/2$. Set $\mathcal{V}_{\beta}=K\cap(\mathcal{P}_{2}(\mathbb{R}^{d})\setminus\mathbb{B}_{\mathcal{W}_{2}}(Q,\beta))$ and $\mathcal{V}_{\beta}^{c}=\mathcal{P}_{2}(\mathbb{R}^{d})\setminus\mathcal{V}_{\beta}$. In summary, it holds that
${\bf P}(\mathcal{V}_{\beta}^{c})\leq\epsilon.$ (19)
Moreover, we can assume that $n$ is large enough such that $\mathcal{W}_{2}(Q_{n},Q)\leq\beta/2$, which implies that
$\mathcal{W}_{2}(Q_{n},P)\geq\mathcal{W}_{2}(P,Q)-\mathcal{W}_{2}(Q_{n},Q)\geq\beta/2,$ (20)
for all $P\in\mathcal{V}_{\beta}$. Next, call
$A_{n}^{2}=\int{\left\|\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{\mathbf{x}-T_{Q_{n},P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q_{n})}\right]\right\|^{2}}\mathrm{d}Q_{n}(\mathbf{x})$
and
$A^{2}=\int{\left\|\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{\mathbf{x}-T_{Q,P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}\right]\right\|^{2}}\mathrm{d}Q(\mathbf{x}).$
The result follows by showing that $A_{n}^{2}\to A^{2}$. Triangle inequality implies that
$\left|A_{n}-\left(\underbrace{\int{\left\|\mathbb{E}_{P\sim\mathbf{P}}\left[{\bf 1}_{\mathcal{V}_{\beta}}(P)\frac{\mathbf{x}-T_{Q_{n},P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q_{n})}\right]\right\|^{2}}\mathrm{d}Q_{n}(\mathbf{x})}_{=:B_{n}^{2}}\right)^{\frac{1}{2}}\right|\\ \leq\left(\int{\left\|\mathbb{E}_{P\sim\mathbf{P}}\left[{\bf 1}_{\mathcal{V}_{\beta}^{c}}(P)\frac{\mathbf{x}-T_{Q_{n},P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q_{n})}\right]\right\|^{2}}\mathrm{d}Q_{n}(\mathbf{x})\right)^{\frac{1}{2}},$
so that, arguing as in Subsection A.1 and using (19), we derive the bound $|A_{n}-B_{n}|\leq\epsilon$ for all $n\in\mathbb{N}$. By the same means $|A-B|\leq\epsilon$ where
$B^{2}=\int{\left\|\mathbb{E}_{P\sim\mathbf{P}}\left[{\bf 1}_{\mathcal{V}_{\beta}}(P)\frac{\mathbf{x}-T_{Q,P}(\mathbf{x})}{\mathcal{W}_{2}(P,Q)}\right]\right\|^{2}}\mathrm{d}Q(\mathbf{x}).$
Therefore, since $\epsilon$ is arbitrary, the result follows after showing that $B_{n}\to B$. To do so, we set $\mathbf{X}_{n}\sim Q_{n}$ for $n\in\mathbb{N}$, $\mathbf{X}\sim Q$ and $P,P^{\prime}\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Arguing as in the proof of Theorem 2.1 in [17] we get for every $P,P^{\prime}\in\mathcal{P}_{2}(\mathbb{R}^{d})$,
[…]
Indeed a straightforward adaptation of the arguments there shows first that there is a limit in distribution which is the distribution of the random vector
$(\mathbf{Z}_{1},\mathbf{Z}_{2},\mathbf{Z}_{3}),$
[…]
$\mathbf{Z}_{2}=T_{Q,P}(\mathbf{Z}_{1}).$
Similarly,
$\mathbf{Z}_{3}=T_{Q,P^{\prime}}(\mathbf{Z}_{1})$
and thus (21) holds. The continuous mapping theorem with the function $(\mathbf{x},\mathbf{y},\mathbf{z})\mapsto(\mathbf{y}-\mathbf{x},\mathbf{z}-\mathbf{x})$ implies that
[…]
Since for all $P\in\mathcal{P}_{2}(\mathbb{R}^{d})$, it holds that $\mathcal{W}_{2}(Q_{n},P)\to\mathcal{W}_{2}(Q,P)$, Slutsky’s theorem yields
[…]
for all $P,P^{\prime}$ such that $\mathcal{W}_{2}(Q,P)>0$ and $\mathcal{W}_{2}(Q,P^{\prime})>0$. As a consequence, (22) holds for ${\bf P}$-a.e. $P,P^{\prime}$.
Let ${\bf P}_{\beta}$ be the probability measure $A\mapsto{\bf P}_{\beta}(A)=\frac{{\bf P}(\mathcal{V}_{\beta}\cap A)}{{\bf P}(\mathcal{V}_{\beta})}$. Therefore, for $(P,P^{\prime})\sim{\bf P}_{\beta}\otimes{\bf P}_{\beta}$ with $({P},P^{\prime})$ independent of $\{\mathbf{X}_{n}\}_{n\in\mathbb{N}}$, we obtain
[…]
Indeed, for a bounded continous function $F:\mathbb{R}\to\mathbb{R}$,
$\mathbb{E}\left[F({\mathbf{Y}}_{n})\right]$ $=\mathbb{E}\left[\mathbb{E}\left[\left.F({\mathbf{Y}}_{n})\right|P,P^{\prime}\right]\right]$
$=\int\int\mathbb{E}\left[\left.F({\mathbf{Y}}_{n})\right|P=\tilde{P},P^{\prime}=\tilde{P}^{\prime}\right]\mathrm{d}{\bf P}_{\beta}(\tilde{P})\mathrm{d}{\bf P}_{\beta}(\tilde{P}^{\prime})$
$=\int\int\mathbb{E}\left[F\left(\frac{\left\langle\mathbf{X}_{n}-T_{Q_{n},\tilde{P}}(\mathbf{X}_{n}),\mathbf{X}_{n}-T_{Q_{n},\tilde{P}^{\prime}}(\mathbf{X}_{n})\right\rangle}{\mathcal{W}_{2}(\tilde{P},Q_{n})\mathcal{W}_{2}(\tilde{P}^{\prime},Q_{n})}\right)\right]\mathrm{d}{\bf P}_{\beta}(\tilde{P})\mathrm{d}{\bf P}_{\beta}(\tilde{P}^{\prime})$
$\underset{n\to\infty}{\longrightarrow}\int\int\mathbb{E}\left[F\left(\frac{\left\langle\mathbf{X}-T_{Q,\tilde{P}}(\mathbf{X}),\mathbf{X}-T_{Q,\tilde{P}^{\prime}}(\mathbf{X})\right\rangle}{\mathcal{W}_{2}(\tilde{P},Q)\mathcal{W}_{2}(\tilde{P}^{\prime},Q)}\right)\right]\mathrm{d}{\bf P}_{\beta}(\tilde{P})\mathrm{d}{\bf P}_{\beta}(\tilde{P}^{\prime})$
$=\mathbb{E}[F(\mathbf{Y})],$
where the above limit holds due to dominated convergence
[…]
$B_{n}^{2}$ $={\bf P}(\mathcal{V}_{\beta})^{2}{\mathbb{E}\left[\frac{\left\langle\mathbf{X}_{n}-T_{Q_{n},P}(\mathbf{X}_{n}),\mathbf{X}_{n}-T_{Q_{n},P^{\prime}}(\mathbf{X}_{n})\right\rangle}{\mathcal{W}_{2}(P,Q_{n})\mathcal{W}_{2}(P^{\prime},Q_{n})}\right]}={\bf P}(\mathcal{V}_{\beta})^{2}\mathbb{E}[\tilde{\mathbf{Y}}_{n}]$
and
$B^{2}$ $={\bf P}(\mathcal{V}_{\beta})^{2}{\mathbb{E}\left[\frac{\left\langle\mathbf{X}-T_{Q,P}(\mathbf{X}),\mathbf{X}-T_{Q,P^{\prime}}(\mathbf{X})\right\rangle}{\mathcal{W}_{2}(P,Q)\mathcal{W}_{2}(P^{\prime},Q)}\right]}={\bf P}(\mathcal{V}_{\beta})^{2}\mathbb{E}[\tilde{\mathbf{Y}}],$
we only need to prove that $\mathbf{Y}_{n}$ is uniformly integrable. The bound (20) implies that it is enough to show that each of the terms of the right hand side of
$|\left\langle\mathbf{X}_{n}-T_{Q_{n},P}(\mathbf{X}_{n}),\mathbf{X}_{n}-T_{Q_{n},P^{\prime}}(\mathbf{X}_{n})\right\rangle|\\ \leq\|\mathbf{X}_{n}\|^{2}+\|T_{Q_{n},P}(\mathbf{X}_{n})\|\|\mathbf{X}_{n}\|+\|T_{Q_{n},P^{\prime}}(\mathbf{X}_{n})\|\|\mathbf{X}_{n}\|+\|T_{Q_{n},P}(\mathbf{X}_{n})\|\|T_{Q_{n},P^{\prime}}(\mathbf{X}_{n})\|$ (23)
are uniformly integrable. Recall that a set $S$ of random variables is uniformly integrable if
$\lim_{R\to+\infty}\sup_{U\in S}\mathbb{E}[|U|{\bf 1}_{|U|>R}]=0.$
Since $\mathcal{V}_{\beta}$ and $\{Q_{n}\}_{n\in\mathbb{N}}$ are relatively compact subsets in the 2-Wasserstein topology, Theorem 7.12 in [42] implies that
$\lim_{R\to+\infty}\sup_{P\in\mathcal{V}_{\beta}}\int_{{\|\mathbf{x}\|^{2}>R}}\|\mathbf{x}\|^{2}{\rm d}P(\mathbf{x})=0$ (24)
and
$\lim_{R\to+\infty}\sup_{n\in\mathbb{N}}\int_{{\|\mathbf{x}\|^{2}>R}}\|\mathbf{x}\|^{2}{\rm d}Q_{n}(\mathbf{x})=0.$ (25)
The last limit (25) implies that the sequence $\{\|\mathbf{X}_{n}\|^{2}\}_{n\in\mathbb{N}}$ is uniformly integrable, so that the first term of the right-hand-side of (23) is uniformly integrable. For the second, we observe that
$\mathbb{E}[$ $\|T_{Q_{n},P}(\mathbf{X}_{n})\|\|\mathbf{X}_{n}\|{\bf 1}_{\|T_{Q_{n},P}(\mathbf{X}_{n})\|\|\mathbf{X}_{n}\|>R}]$
$\leq\mathbb{E}\left[\|T_{Q_{n},P}(\mathbf{X}_{n})\|\|\mathbf{X}_{n}\|{\bf 1}_{\|\mathbf{X}_{n}\|>R^{\frac{1}{2}}}\right]+\mathbb{E}\left[\|T_{Q_{n},P}(\mathbf{X}_{n})\|\|\mathbf{X}_{n}\|{\bf 1}_{\|T_{Q_{n},P}(\mathbf{X}_{n})\|>R^{\frac{1}{2}}}\right]$
$\leq\left(\mathbb{E}\left[\|T_{Q_{n},P}(\mathbf{X}_{n})\|^{2}\right]\mathbb{E}\left[\|\mathbf{X}_{n}\|^{2}{\bf 1}_{\|\mathbf{X}_{n}\|>R^{\frac{1}{2}}}\right]\right)^{\frac{1}{2}}$
$\qquad\qquad+\left(\mathbb{E}\left[\|T_{Q_{n},P}(\mathbf{X}_{n})\|{\bf 1}_{\|T_{Q_{n},P}(\mathbf{X}_{n})\|>R^{\frac{1}{2}}}\right]\mathbb{E}\left[\|\mathbf{X}_{n}\|^{2}\right]\right)^{\frac{1}{2}}$
$\leq\left(\sup_{P\in\mathcal{V}_{\beta}}\int\|\mathbf{x}\|^{2}{\rm d}P(\mathbf{x})\int_{\|\mathbf{x}\|^{2}>R}\|\mathbf{x}\|^{2}{\rm d}Q_{n}(\mathbf{x})\right)^{\frac{1}{2}}$
$\qquad\qquad+\left(\sup_{P\in\mathcal{V}_{\beta}}\int_{\|\mathbf{x}\|^{2}>R}\|\mathbf{x}\|^{2}{\rm d}P(\mathbf{x})\int\|\mathbf{x}\|^{2}{\rm d}Q_{n}(\mathbf{x})\right)^{\frac{1}{2}},$
where we used the fact that $T_{Q_{n},P}(\mathbf{X}_{n})\sim P$ for all $n\in\mathbb{N}$. Since, $\sup_{n\in\mathbb{N}}\int\|\mathbf{x}\|^{2}{\rm d}Q_{n}(\mathbf{x})$ and $\sup_{P\in\mathcal{V}_{\beta}}\int\|\mathbf{x}\|^{2}{\rm d}P(\mathbf{x})$ are bounded, the previous display, (24) and (25) imply that the second term of (23) is uniformly integrable. Since $P$ and $P^{\prime}$ are exchangeable, the same holds for the third term. The uniform integrability of the last one follows directly from (24).
Proof of Lemma 5.5.
From [43, Corollary 5.23], for every $\epsilon>0$, it holds that $Q(|T_{Q,P_{n}}-T_{Q,P}|\geq\epsilon)\to 0$. As $\|T_{Q,P_{n}}-T_{Q,P}\|_{L^{2}(Q)}$ is uniformly bounded, the sequence $\{T_{Q,P_{n}}-T_{Q,P}\}_{n\in\mathbb{N}}$ is compact w.r.t. the weak topology of $L^{2}(Q)$ by the Banach-Alaoglu–Bourbaki theorem (cf. [8, Theorem 3.16]). Therefore, for each subsequence $\{T_{Q,P_{n_{k}}}-T_{Q,P}\}_{k\in\mathbb{N}}$ there exists a further subsequence $\{T_{Q,P_{n_{k_{\ell}}}}-T_{Q,P}\}_{\ell\in\mathbb{N}}$ such that
$\langle T_{Q,P_{n_{k_{\ell}}}}-T_{Q,P},h\rangle_{L^{2}(Q)}\to\langle L,h\rangle_{L^{2}(Q)}$
for some $L\in L^{2}(Q)$ and all $h\in L^{2}(Q)$. We prove now that $L=0$, irrespective of the subsequences. To improve readability, we write $\{T_{Q,P_{n}}-T_{Q,P}\}_{n\in\mathbb{N}}$ instead of $\{T_{Q,P_{n_{k_{\ell}}}}-T_{Q,P}\}_{\ell\in\mathbb{N}}$. Since $Q(|T_{Q,P_{n}}-T_{Q,P}|\geq\epsilon)\to 0$ and
$\|T_{Q,P_{n}}\|_{L^{2}(Q)}^{2}=\int\|\mathbf{x}\|^{2}{\rm d}P_{n}(\mathbf{x})\to\int\|\mathbf{x}\|^{2}{\rm d}P(\mathbf{x})=\|T_{Q,P}\|^{2}_{L^{2}(Q)}<+\infty,$
Vitali convergence theorem implies that $\{T_{Q,P_{n}}-T_{Q,P}\}_{n\in\mathbb{N}}$ converges to zero in the reflexive space $L^{\frac{3}{2}}(Q)$. Therefore, 00 is also the weak limit of $T_{Q,P_{n}}-T_{Q,P}$ in $L^{\frac{3}{2}}(Q)$, i.e.,
[…]
for all $h\in L^{3}(Q)$. As a consequence, $L=0$, $Q$-a.e. Moreover,
$\|T_{Q,P_{n}}-T_{Q,P}\|_{L^{2}(Q)}^{2}$ $=\|T_{Q,P_{n}}\|_{L^{2}(Q)}^{2}+\|T_{Q,P}\|_{L^{2}(Q)}^{2}-2\langle T_{Q,P_{n}},T_{Q,P}\rangle_{L^{2}(Q)}$
$\to 2\|T_{Q,P}\|^{2}-2\langle T_{Q,P},T_{Q,P}\rangle_{L^{2}(Q)}$
$=0.$
Proof of Theorem 5.6.
Fix $\epsilon>0$. As $\mathbf{P}\in\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ is atomless there exists an open Wasserstein ball
$\mathbb{B}_{\mathcal{W}_{2}}(Q,\beta)=\{P\in\mathcal{P}_{2}(\mathbb{R}^{d}):\ \mathcal{W}_{2}(P,Q)<\beta\}$
[…]
${\bf P}_{n}(\mathbb{B}_{\mathcal{W}_{2}}(Q,\beta/2))\leq\epsilon/4\quad\text{for all }n\geq n_{0}.$
As $\{{\bf P}_{n}\}_{n\in\mathbb{N}}\subset\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ is tight, there exists a compact set $K\subset\mathcal{P}_{2}(\mathbb{R}^{d})$ such that
${\bf P}_{n}(\mathcal{P}_{2}(\mathbb{R}^{d})\setminus K)\leq\epsilon/4\quad\text{for all }n\geq n_{0}.$
Call $V=K\cap(\mathcal{P}_{2}(\mathbb{R}^{d})\setminus\mathbb{B}_{\mathcal{W}_{2}}(Q,\beta/2))$ and $V^{c}=\mathcal{P}_{2}(\mathbb{R}^{d})\setminus V$. Then
${\bf P}(V^{c})+{\bf P}_{n}(V^{c})\leq\epsilon\quad\text{for all }n\geq n_{0}.$
We call
$A:=\left|\left\|\int\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}\mathrm{d}{\bf P}_{n}(P)\right\|_{L^{2}(Q)}-\left\|\int\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}\mathrm{d}{\bf P}(P)\right\|_{L^{2}(Q)}\right|.$
The triangle inequality yields
$A$ $\leq\left\|\int\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}\mathrm{d}({\bf P}_{n}-{\bf P})(P)\right\|_{L^{2}(Q)}$
$\leq\left\|\int_{V}\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}\mathrm{d}({\bf P}_{n}-{\bf P})(P)\right\|_{L^{2}(Q)}+\left\|\int_{V^{c}}\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}\mathrm{d}{\bf P}(P)\right\|_{L^{2}(Q)}$
$\qquad\qquad+\left\|\int_{V^{c}}\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}\mathrm{d}{\bf P}_{n}(P)\right\|_{L^{2}(Q)}.$
Arguing as in Section A.1 we get that
$\left\|\int_{V^{c}}\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}\mathrm{d}{\bf P}_{n}(P)\right\|_{L^{2}(Q)}+\left\|\int_{V^{c}}\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}\mathrm{d}{\bf P}_{n}(P)\right\|_{L^{2}(Q)}\leq{\bf P}(V^{c})+{\bf P}_{n}(V^{c})\leq\epsilon.$
Moreover, as the function
$V\ni P\mapsto\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}\in L^{2}(Q)$
is continuous and bounded (Lemma 5.5), for every $h\in L^{2}(Q)$ it holds that
$\int_{V}\left\langle\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)},h\right\rangle_{L^{2}(Q)}\mathrm{d}({\bf P}_{n}-{\bf P})(P)\to 0,$
meaning that $\int_{V}\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}\mathrm{d}({\bf P}_{n}-{\bf P})(P)$ converges to zero in the weak topology of $L^{2}(Q)$. However, as the set
$\left\{\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}:P\in V\right\}\cup\{0\}$
is compact (note that $V$ is compact in $\mathcal{P}_{2}(\mathbb{R}^{d})$ and $\mathcal{P}_{2}(\mathbb{R}^{d})\backslash\{Q\}\ni P\mapsto\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}$ is continuous, see Lemma 5.5), its closed convex hull, namely $C$, is compact as well. Since $\int_{V}\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}\mathrm{d}{\bf P}_{n}$ lies in $C$ for all $n\in\mathbb{N}$, the convergence of $\int_{V}\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}\mathrm{d}({\bf P}_{n}-{\bf P})(P)$ towards zero holds in the strong topology of $L^{2}(Q)$. We have proven that $A\leq\epsilon$ for $n$ big enough. Since $\epsilon$ was arbitrarily chosen, the result follows. ∎
Appendix D Proof of Lemma 6.2
Let $S\subset\mathcal{P}_{p}(\mathbb{R}^{d})$ be a closed set and define
${\rm BL}_{1}(S)=\left\{f:S\to\mathbb{R}:\ |f(P)|\leq 1\ {\rm and}\ |f(P)-f(Q)|\leq\mathcal{W}_{p}(P,Q),\ \forall\,P,Q\in S\right\}.$
Fix $f\in{\rm BL}_{1}(\mathcal{P}_{p}(\mathbb{R}^{d}))$. Then
$\left|\int f(P)\mathrm{d}({\bf P}_{n,m}-{\bf P})(P)\right|\leq\underbrace{\left|\int f(P)\mathrm{d}({\bf P}_{n,m}-{\bf P}_{n})(P)\right|}_{A_{n,m}(f)}+\underbrace{\left|\int f(P)\mathrm{d}({\bf P}_{n}-{\bf P})(P)\right|}_{B_{n}(f)},$
where ${\bf P}_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{P_{i}}$. It can be proved by standard means that
$\mathbb{E}\left[\sup_{f\in{\rm BL}_{1}(\mathcal{P}_{p}(\mathbb{R}^{d}))}B_{n}(f)\right]\to 0$
as $n\to\infty$. Since $f\in{\rm BL}_{1}(\mathcal{P}_{p}(\mathbb{R}^{d}))$, it holds that
$A_{n,m}(f)=\left|\frac{1}{n}\sum_{i=1}^{n}f(P_{i,m})-f(P_{i})\right|\leq\frac{1}{n}\sum_{i=1}^{n}\min(2,\mathcal{W}_{p}(P_{i,m},P_{i}))$
which, by taking expectations, implies
$\mathbb{E}\left[\sup_{f\in{\rm BL}_{1}(\mathcal{P}_{p}(\mathbb{R}^{d}))}A_{n,m}(f)\right]\leq\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\min(2,\mathcal{W}_{p}(P_{i,m},P_{i}))].$
Since the sequence $\{\mathcal{W}_{p}(P_{i,m},P_{i})\}_{i=1}^{n}$ is exchangeable, it holds that
$\mathbb{E}\left[\sup_{f\in{\rm BL}_{1}(\mathcal{P}_{p}(\mathbb{R}^{d}))}A_{n,m}(f)\right]\leq\mathbb{E}[\min(2,\mathcal{W}_{p}(P_{1,m},P_{1}))].$
The latter tends to zero by Glivenko–Cantelli theorem and the fact that
$\frac{1}{m}\sum_{j=1}^{m}\mathbf{X}_{1,j}^{p}\xrightarrow{a.s.}\int\|\mathbf{x}\|^{p}\mathrm{d}P_{1}(\mathbf{x})$
as $m\to\infty$.